### v5

# The Hasse Principle for Geometric Variational Problems: An Illustration via Area-minimizing Submanifolds

Abstract.
The Hasse principle in number theory states that information about integral solutions to Diophantine equations can be pieced together from real solutions and solutions modulo prime powers. We show that the Hasse principle holds for area-minimizing submanifolds: information about area-minimizing submanifolds in integral homology can be fully recovered from those in real homology and mod $n$ homology for all $n\in\mathbb{Z}_{\geq 2}$. As a consequence we derive several surprising conclusions, including: area-minimizing submanifolds in mod $n$ homology are asymptotically much smoother than expected, area-minimizing submanifolds are not generically calibrated, and products of area-minimizing submanifolds are not generically area-minimizing. We conjecture that the Hasse principle holds for all geometric variational problems that can be formulated on chain spaces over different coeffiicients, e.g., Almgren-Pitts min-max, mean curvature flow, Song’s spherical Plateau problem, minimizers of elliptic and other general functionals, etc.
1. Introduction
The Hasse principle [19, 44] in number theory states that information about integral solutions to Diophantine equations can be pieced together from real solutions and all solutions modulo prime powers. An illustrative example is the Hasse-Minkowski theorem for quadratic forms [43], and its vast generalization Siegel’s weight formula [12], the latter which of uses the number of solutions modulo prime powers to give a weighted sums of number of solutions over $\mathbb{Z}$. Another example is the original formulation of the Birch and Swinnerton-Dyer conjecture [3], where the number of solutions modulo primes conjecturally gives ranks of rational solutions of elliptic curves. Though the Hasse principle does not hold in general [42], the guiding philosophy of “local-to-global” reasoning remains a cornerstone of modern number theory.
A striking analytical analogue lies dormant in Riemannian geometry and geometric measure theory. The classical variational problem of finding area-minimizing representatives in integral homology classes is solved by Federer and Fleming ([15]):
Every integral homology class on a compact Riemannian manifold admits an area-minimizing representative.
Their minimizers live in the space of integral currents, i.e., the Whitney-flat closure of integer-coefficient polyhedral chains [15]. Parallel existence theorems hold for real homology and for homology with $\mathbb{Z}/n\mathbb{Z}$ coefficients for all $n\in\mathbb{Z}_{\geq 2}$, established respectively in [14] and [16].
According to Ulam [49], Banach often remarked
Good mathematicians see analogies between theorems or theories; the very best ones see analogies between analogies.
Inspired by this maxim, we explore how the local-to-global insight of the Hasse principle resonates with the theory for area-minimizing currents, two problems that, at first glance, live worlds apart. We prove the following Hasse principle for area-minimizing currents:
Theorem 1.
Area-minimizing submanifolds in integral homology can be fully recovered from those in real homology and mod $n$ homology for all $n\in\mathbb{Z}_{\geq 2}.$
To state our result more precisely, let us start with some basic assumptions and definitions that we will use throughout this manuscript.
Assumption 1.0.1.
Assume that
  • $M$ is a $(d+c)$-dimensional compact closed not necessarily orientable smooth manifold, with $d,c\in\mathbb{Z}_{\geq 0},$
  • $\Sigma]\in H_{d}(M,\mathbb{Z})$ is a $d$-dimensional integral homology class,
  • $g$ is a smooth Riemannian metric on $M.$
Definition 1.0.1.
We classify the integral homology class $\Sigma]$ according to its algebraic properties.
  • We say $\Sigma]$ is a torsion class if there is an integer $v$ such that $v[\Sigma]=0.$ Otherwise we say $\Sigma]$ is a free class,
  • If $\Sigma]$ is a free class, $M$ is orientable, and $d\geq 1,c\geq 2$, then the representable multiplicity $\mu_{[\Sigma]}$ of $\Sigma]$ is defined to be the smallest integer $\mu_{[\Sigma]}\in\mathbb{Z}_{\geq 2}$ such that $\mu_{[\Sigma]}[\Sigma]$ can be represented by a smoothly embedded connected submanifold.
The well-definedness of $\mu_{[\Sigma]}$ is proved in Lemma 2.1.2. Recall [20] that $H_{d}(M,\mathbb{Z})$ is a finitely generated abelian group which can be written as a splitting up to choices of basis,
(1.0.1) $H_{d}(M,\mathbb{Z})=\mathbb{Z}^{b_{d}}\oplus\bigoplus_{i\in I}\mathbb{Z}/p_{i}^{\nu_{i}}\mathbb{Z},$
where $b_{d}$ is the $d$-th Betti number of $M$, II is a finite set of indices, the $p_{i}$ are not necessarily distinct prime numbers and $\nu_{i}\in\mathbb{Z}_{\geq 1}$. By [29, Theorem 1.57(c)], the set $\{b_{d}\}\cup_{i\in I}\{p_{i},v_{i}\}$ is uniquely determined by $M.$
Definition 1.0.2.
Define the $d$-th torsion number $\tau_{d}$ of $M$ to be
$\tau_{d}=\prod_{i\in I}p_{i}^{\nu_{i}}.$
Set $\tau_{d}=1$ if $H_{d}(M,\mathbb{Z})$ is torsion-free.
Like the $d$-th Betti number $b_{d}$, the $d$-th torsion number $\tau_{d}$ is an invariant of $M.$
By the universal coefficient theorem [20, Theorem 3A.3], for any coefficient ring $R$, we have a natural short exact sequence
(1.0.2) $0\rightarrow H_{d}(M,\mathbb{Z})\otimes R\overset{\iota}{\rightarrow}H_{d}(M,R)\rightarrow\operatorname{Tor}(H_{d-1}(M,\mathbb{Z}),R)\to 0.$
Definition 1.0.3.
The $R$-reduction $\Sigma]^{R}$ of the integral homology class $\Sigma]$ is defined to be
$[\Sigma]^{R}=\iota([\Sigma]\otimes 1)\in H_{d}(M,R),$
where $\iota$ is the natural injection in (1.0.2) and $1$ is the unit element in $R.$
We need a norm $\left\lVert\cdot\right\rVert^{R}$ on $R$ in order to measure the area of chains in $R$ homology.
Definition 1.0.4.
We say a map $\left\lVert\cdot\right\rVert^{R}:R\to\mathbb{R}_{\geq 0}$ is a norm on $R$ if $\left\lVert\cdot\right\rVert^{R}$ satisfies the following three properties:
  • (reversibility) for all $e\in R$, we have
    $\left\lVert-e\right\rVert^{R}=\left\lVert e\right\rVert^{R},$
  • (triangle inequality) for all $e,f\in R$, we have
    $\left\lVert e+f\right\rVert^{R}\leq\left\lVert e\right\rVert^{R}+\left\lVert f\right\rVert^{R}.$
  • (positivity) we have the propositional equivalence, for all $e\in R,$
    $\left\lVert e\right\rVert^{R}=0\Leftrightarrow e=0.$
We want to emphasize that we do not require the usual homogeneity axiom, i.e., $\left\lVert ve\right\rVert^{R}=|v|\left\lVert e\right\rVert^{R}$ for $e\in R,v\in\mathbb{Z}$ on $R,$ but only use the weaker reversibility axiom. This is an intentional omission in order to measure elements in groups with torsions.
Recall [16, 53, 55] that $\left\lVert\cdot\right\rVert^{R}$ induces a natural flat distance $\mathcal{F}^{R}$ and mass norm $\mathbf{M}^{R}_{g}$ on the space of $R$-coefficient polyhedron chains on the Riemannian manifold $(M,g)$.
Definition 1.0.5.
The $R$-flat chain $\mathcal{F}_{g}(M,R)$ in the metric $g$ is the $\mathcal{F}^{R}$ closure of the set of $R$-coefficient polyhedron chains.
Intuitively speaking the mass $\mathbf{M}_{g}^{R}$ of a $d$-dimensional chain is just the $d$-dimensional area of a chain counted with multiplicity by $\left\lVert\cdot\right\rVert^{R}.$ For instance, by Definition 1.0.8,
$\mathbf{M}^{\mathbb{Z}/4\mathbb{Z}}(\pm 2S^{1})=\left\lVert\pm 2\right\rVert^{\mathbb{Z}/4\mathbb{Z}}\mathbf{M}^{\mathbb{Z}/4\mathbb{Z}}(S^{1})=2\cdot 2\pi.$
By [13, Section 4.2.16], the finite area $\mathbb{Z}$-coefficient flat chains representing $\Sigma]\in H_{d}(M,\mathbb{Z})$ are precisely the classical integral currents representing $\Sigma]$. As we are concerned with area-minimization, we will use the literature convention of using integral currents to represent integral homology.
Definition 1.0.6.
When the canonical homomorphism $\mathbb{Z}\to R$ is norm non-increasing, define the $R$-reduction $T^{R}$ of an integral current $T\in[\Sigma]$ to be
$T^{R}=\iota(T\otimes 1)\in[\Sigma]^{R}.$
Here $\iota$ is the canonical homomorphism in (1.0.2), which also exists on the chain level. It is straightforward to verify that $T^{R}$ is well-defined.
We want to emphasize that the set of $R$-flat chains in $\Sigma]^{R}$ is much larger than the $R$-reduction of integral currents in $\Sigma]$. For instance, in $S^{4}$, where the second homology is trivial, an embedded $\mathbb{RP}^{2}$ is a mod $2$ cycle, but is not a $\mathbb{Z}/2\mathbb{Z}$-reduction of any integral cycle.
Recall Definition 1.0.5.
Definition 1.0.7.
Define the $R$-area $\left\lVert[\Sigma]\right\rVert^{R}_{g}$ of $\Sigma]^{R}$ to be
(1.0.3) $\left\lVert[\Sigma]\right\rVert^{R}_{g}=\inf_{\begin{subarray}{c}T\in\mathcal{F}_{g}(M,R)\\ T\in[\Sigma]^{R}\end{subarray}}\mathbf{M}_{g}^{R}(T).$
We say an $R$-flat chain $T\in[\Sigma]^{R}$ is $R$-area-minimizing in $\Sigma]^{R}$ with respect to metric $g,$ if $T$ achieves the $R$-area of $\Sigma]^{R}$ defined in (1.0.3), i.e.,
$\mathbf{M}^{R}_{g}(T)=\left\lVert[\Sigma]\right\rVert^{R}_{g}.$
Define the set of $R$-area-minimizing representatives $\mathbf{MIN}_{g}^{R}\left([\Sigma]\right)$ of $\Sigma]^{R}$ to be
$\mathbf{MIN}_{g}^{R}\left([\Sigma]\right)=\{T|T\in\mathcal{F}_{g}(M,R),T\in[\Sigma]^{R},\mathbf{M}^{R}_{g}(T)=\left\lVert[\Sigma]\right\rVert^{R}_{g}\}.$
Intuitively speaking, $\left\lVert[\Sigma]\right\rVert^{R}_{g}$ is the infimum of area of all $R$-coefficient polyhedron chains representing the $R$-reduction $\Sigma]^{R}$ of $\Sigma]$. By the discussion before Definition 1.0.7, we have
Fact 1.0.1.
The following inequality holds
(1.0.4) $\left\lVert[\Sigma]\right\rVert^{R}_{g}\leq\inf_{T\in[\Sigma]}\mathbf{M}_{g}^{R}(T^{R}).$
In this manuscript, we will always use the following classical norm conventions:
Definition 1.0.8.
Define
$\left\lVert\cdot\right\rVert^{\mathbb{Z}}=$ $|\cdot|,$
$\left\lVert\cdot\right\rVert^{\mathbb{R}}=$ $|\cdot|,$
$\left\lVert\cdot\right\rVert^{\mathbb{Z}/n\mathbb{Z}}=$ $|\cdot|_{(-\frac{n}{2},\frac{n}{2}]}.$
Here $|\cdot|$ is the absolute value on $\mathbb{R}$ and $|\cdot|_{(-\frac{n}{2},\frac{n}{2}]}$ means the absolute value of the unique lift of an element in $\mathbb{Z}/n\mathbb{Z}$ to $(-\frac{n}{2},\frac{n}{2}]$. With $\left\lVert\cdot\right\rVert^{R}$ defined as above, when $R=\mathbb{Z},\mathbb{R}$ or $\mathbb{Z}/n\mathbb{Z}$ with $n\in\mathbb{Z}_{\geq 2},$ there always exists at least one area-minimizing $R$-flat chain representative of $\Sigma]^{R} [$14, 16, 13].
The following assumption will always be assumed.
Assumption 1.0.2.
Set $R=\mathbb{R}$ or $\mathbb{Z}$ or $\mathbb{Z}/n\mathbb{Z}$ for $n\in\mathbb{Z}_{\geq 2}$ and use Definition 1.0.8 for $\left\lVert\cdot\right\rVert^{R}.$
It is straightforward to verify that by Definition 1.0.8, the canonical ring homomorphism $\mathbb{Z}\to R$ is norm non-increasing. Thus, we always have
$\mathbf{M}^{R}_{g}(T^{R})\leq\mathbf{M}_{g}^{\mathbb{Z}}(T).$
By (1.0.4), we deduce that
Fact 1.0.2.
The following inequality always holds
(1.0.5) $\left\lVert[\Sigma]\right\rVert^{R}_{g}\leq\left\lVert[\Sigma]\right\rVert^{\mathbb{Z}}_{g}.$
We emphasize again that $R$-reduction of a homology class introduces many more representatives, e.g., unorientable ones for $R=\mathbb{Z}/2\mathbb{Z}$, $Y$-shaped union of three rays starting at the origin for $R=\mathbb{Z}/3\mathbb{Z},$ the operator of $L^{2}$ inner product with coclosed differential forms for $R=\mathbb{R},$ etc. Thus, it is not expected and in fact not true in general that equality in (1.0.5) can hold. When (1.0.5) is strict inequality, we may even have
(1.0.6) $\mathbf{MIN}_{g}^{R}\left([\Sigma]\right)\cap\left(\mathbf{MIN}_{g}^{\mathbb{Z}}\left([\Sigma]\right)\right)^{R}=\emptyset.$
Now we are ready to state our more detailed version of Theorem 1. Recall Assumption 1.0.1 and Definition 1.0.2.
Theorem 2.
For all $\Sigma]\in H_{d}(M,\mathbb{Z})$, there exists a positive integer $N_{[\Sigma],g}$, and an open set $\Omega_{[\Sigma],g}$ containing $g$ in the space of Riemannian metrics, both depending only on $\Sigma]$ and $g$, such that for all metrics $h\in\Omega_{[\Sigma],g},$ all $n\geq N_{[\Sigma],g}$ with $\tau_{d}|n$,
  • equality holds in (1.0.5), i.e., $\left\lVert[\Sigma]\right\rVert_{h}^{\mathbb{Z}/n\mathbb{Z}}=\left\lVert[\Sigma]\right\rVert_{h}^{\mathbb{Z}},$
  • the $\mathbb{Z}/n\mathbb{Z}$-reduction restricted to $\mathbf{MIN}_{g}^{\mathbb{Z}}\left([\Sigma]\right)$ is a one to one map from   $\mathbf{MIN}_{g}^{\mathbb{Z}}\left([\Sigma]\right)$ to $\mathbf{MIN}_{g}^{\mathbb{Z}/n\mathbb{Z}}\left([\Sigma]\right)$.
Here the inverse to $\mathbb{Z}/n\mathbb{Z}$-reduction restricted to $\mathbf{MIN}_{g}^{\mathbb{Z}}\left([\Sigma]\right)$-reduction equals Federer’s ”representative” modulo $n$ [13, Section 4.2.26]. Unfortunately, Federer used the word ”representative” in a totally different way than we did, and the reader should not confuse Federer’s ”representative” with our use of representatives of homology classes.
Though the statement of Theorem 2 only describes information from $\mathbb{Z}/n\mathbb{Z}$ homologies, like the proof of the Hasse-Minkowski theorem, information obtained from both real homology and $\mathbb{Z}/n\mathbb{Z}$ homologies is indispensable to the proof of Theorem 2 .
State of art regularity proved in [45, 8, 9, 7, 6] by De Lellis and his collaborators shows that the singular set of $T$ is of Hausdorff dimension
[…]
See also [24, 23, 22] for an independent proof of the first case. There is a $1$-dimension gap between these two results. In the case of $c=1,$ the singular sets of area-minimizing representatives of integral homology classes have Hausdorff dimension at most (d7)(d-7) [39], and the gap is increased to $6$ dimensions in general. The dimension gap in regularity cannot be overcome, as the regularity results above are sharp. However, as a corollary of Theorem 2, we can prove that
Corollary 1.0.1.
Area-minimizing representatives of $\Sigma]^{\mathbb{Z}/n\mathbb{Z}}$ are area-minimizing integral currents for $n$ large enough, thus enjoying much better regularity than general $\mathbb{Z}/n\mathbb{Z}$ area-minimizing representatives.
To the author’s knowledge, the only precursor to the above corollary in the literature is by Frank Morgan [32] for hypersurface solutions to $\mathbb{Z}/n\mathbb{Z}$-Plateau problems with boundaries of at most $n/2$ components on convex sets. Experts view that result as a “special case” [30]. Hence, the universal asymptotic regularity improvement proved here is both unexpected and striking.
One can ask what information about area-minimizing representatives in integral homology is lost in $\mathbb{R}$ and $\mathbb{Z}/n\mathbb{Z}$ homology for any finite sets of $n$. To this end, we have the following theorem, which shows the sharpness of Theorem 2 and the impossibility of finding a uniform upper bound for $N_{[\Sigma],g}$ in Theorem 2. Thus, to recover the $\mathbb{Z}$-area-minimizing representative of $\Sigma]$, we do need information from real homology and all $\mathbb{Z}/n\mathbb{Z}$ homologies, much like the role of real and all mod prime power solutions in the Hasse-Minkowski theorem. Recall Definition 1.0.1.
Theorem 3.
If $d\geq 1,c\geq 2,$ $M$ is orientable, and $\Sigma]$ a free class, then for any real number $\rho>1$ and all integers $n\geq 2$ with $\operatorname{gcd}\left(n,\mu_{[\Sigma]}\right)=1,$  there exists a non-empty open set $\Omega^{\prime}_{[\Sigma],n,\rho}$ in the space of Riemannian metrics, depending on $\Sigma],n,$ and $\rho,$ such that for all metrics $h\in\Omega^{\prime}_{[\Sigma],n,\rho}$ we have
(1.0.7) ${\left\lVert[\Sigma]\right\rVert^{\mathbb{Z}}_{h}}>$ $\rho{\left\lVert[\Sigma]\right\rVert^{\mathbb{Z}/n\mathbb{Z}}_{h}},$
(1.0.8) ${\left\lVert[\Sigma]\right\rVert^{\mathbb{Z}}_{h}}>$ $\rho{\left\lVert[\Sigma]\right\rVert^{\mathbb{R}}_{h}}.$
We want to remark that if $\operatorname{gcd}(\tau_{d},\mu_{[\Sigma]})=1,$ e.g., $\tau_{d}=1,$ then we can always find  $\operatorname{gcd}\left(n,\mu_{[\Sigma]}\right)=1$ and $\tau_{d}|n,$ so that Theorem 3 genuinely shows the sharpness of Theorem 2.
As a byproduct, we prove that,
Corollary 1.0.2.
If $d\geq 1,c\geq 2$, $\mathbb{Z}$-area-minimizing submanifolds on orientable manifolds are not generically calibrated.
In [14], Federer proved that area-minimizing hypersurfaces on orientable manifolds are always calibrated. To the best of the author’s knowledge, all known examples of non-torsion $\mathbb{Z}$-area-minimizing integral currents are calibrated. Thus, Corollary 1.0.2 is both sharp and genuinely unexpected.
When the author was an undergraduate at Duke University, Professor William Allard recalled a conversation he once had with Professor Herbert Federer in which they discussed whether the title, Homological Integration Theory, of the fourth chapter in the draft of the monumental work [13], might be too grand a name for its actual contents. Federer remarked that, in his view, homology constituted the most essential component of the entire theory.
Given Federer’s perspective, we propose an alternative interpretation of our main theorems:
We are quantifying the effect of $R$-reduction in the universal coefficient theorem on area-minimizing representatives.
Theorem 2 says that for integers $n$ large enough and divisible by $\tau_{d}$, the $\mathbb{Z}/n\mathbb{Z}$-reduction sends $\mathbb{Z}$-area-minimizing representatives bijectively to $\mathbb{Z}/n\mathbb{Z}$-area-minimizing representatives, while Theorem 3 shows that in non-empty open sets of metrics $R$-reduction fails drastically at sending $\mathbb{Z}$-area-minimizing representatives to $R$-area-minimizing representatives for $R=\mathbb{R}$ and $R=\mathbb{Z}/n\mathbb{Z}$ with $n\in\mathbb{Z}_{\geq 2}.$
Thus the following problem provides a natural framework for our results.
Problem 1.
For any natural homomorphism among homology groups, quantify the effect of the homomorphism on area-minimizing representatives.
To give another illustration of the above problem beyond our setting of the universal coefficient theorem, we will quantify the effect of Künneth formula on the homology of products. First, let us introduce another manifold.
Assumption 1.0.3.
Assumption 1.0.1 holds with another set of objects $M^{\prime},d^{\prime},c^{\prime},[\Sigma^{\prime}],g^{\prime}$ in place of $M,d,c,[\Sigma],g$.
By the the Künneth formula [20, Theorem 3B.6], for every natural number $l\in\mathbb{N},$ we have,
(1.0.9) $0\to\bigoplus_{i\in\mathbb{N}}H_{i}(M,\mathbb{Z})\otimes H_{l-i}(M^{\prime},\mathbb{Z})\overset{\mathfrak{i}}{\rightarrow}H_{l}(M\times M^{\prime},\mathbb{Z})\to\bigoplus_{i\in\mathbb{N}}\operatorname{Tor}\left(H_{i}(M,\mathbb{Z}),H_{l-i-1}(M^{\prime},\mathbb{Z})\right)\to 0,$
and when $R$ is a field [20, Corollary 3B.7], e.g., $R=\mathbb{R}$ or $R=\mathbb{Z}/p\mathbb{Z}$ with $p$ a prime number,
(1.0.10) $0\to\bigoplus_{i\in\mathbb{N}}H_{i}(M,R)\otimes H_{l-i}(M^{\prime},R)\overset{\mathfrak{i}}{\rightarrow}H_{l}(M\times M^{\prime},R)\to 0.$
Definition 1.0.9.
Define
(1.0.11) $[\Sigma]\times[\Sigma^{\prime}]=\mathfrak{i}\left([\Sigma]\otimes[\Sigma^{\prime}]\right).$
On the chain level, the tensor product on the left hand side of (1.0.9) and (1.0.10) are induced by cross products of chains [20, Section 3.B]. Thus, the product defined in (1.0.11) corresponds to the Cartesian product of $R$-flat chains for $R=\mathbb{Z},\mathbb{R}$ or $\mathbb{Z}/p\mathbb{Z}$ with $p$ a prime number.
As there are many representatives of $\Sigma]\times[\Sigma^{\prime}]$ which do not come from Cartesian products of cycles, we have,
Fact 1.0.3.
The following inequality holds,
(1.0.12) $\left\lVert[\Sigma]\right\rVert^{R}_{g}\times\left\lVert[\Sigma^{\prime}]\right\rVert^{R}_{g^{\prime}}\geq\left\lVert[\Sigma]\times[\Sigma^{\prime}]\right\rVert^{R}_{g\times g^{\prime}}.$
In the positive direction, equality in (1.0.12) holds in many well-known cases. When $R=\mathbb{R},$ by [38, Theorem 5.2], we have,
Fact 1.0.4.
If $d\leq 2,$ or $c\leq 2,$ or both $\Sigma],[\Sigma^{\prime}]$ are of dimension $3$ or codimension $3,$ then we have $\left\lVert[\Sigma]\right\rVert^{\mathbb{R}}_{g}\times\left\lVert[\Sigma^{\prime}]\right\rVert^{\mathbb{R}}_{g^{\prime}}=\left\lVert[\Sigma]\times[\Sigma^{\prime}]\right\rVert^{\mathbb{R}}_{g\times g^{\prime}},$
It is an open problem whether the above fact holds unconditionally for all dimensions and codimensions. See [61] for recent progress.
Federer [14] proved that codimension one $\mathbb{Z}$-area-minimizing representatives on orientable ambient manifolds are also $\mathbb{R}$-area-minimizing representatives. Thus, by Fact 1.0.4, we have that,
Fact 1.0.5.
When both $\Sigma],[\Sigma^{\prime}]$ are of codimension $1$ and both $M,M^{\prime}$ are orientable,
(1.0.13) $\left\lVert[\Sigma]\right\rVert^{\mathbb{Z}}_{g}\times\left\lVert[\Sigma^{\prime}]\right\rVert^{\mathbb{Z}}_{g^{\prime}}=\left\lVert[\Sigma]\times[\Sigma^{\prime}]\right\rVert^{\mathbb{Z}}_{g\times g^{\prime}}.$
In other words, the Cartesian product of $\mathbb{Z}$-area-minimizing hypersurfaces on orientable manifolds are $\mathbb{Z}$-area-minimizing on the Riemannian product of ambient manifolds. Tracing through the proof, one can show that Fact 1.0.5 holds for successive products of more than two homology classes as well.
Examples like products of Kähler subvarieties in products of Kähler manifolds [18] show that (1.0.13) hold in other codimensions in natural settings. By [35] and [26, Lemma 2.11], holomorphic subvarieties of complex flat tori show that (1.0.13) holds with $\mathbb{Z}$ replaced by $\mathbb{Z}/n\mathbb{Z}$ for $n\in\mathbb{Z}_{\geq 2}$ in natural settings as well.
On the other hand, for Plateau problems in Euclidean space, [34, 56] provide examples where Cartesian products of solutions are not solutions to the product boundary conditions. As another byproduct of Theorem 3, we prove that products of area-minimizing submanifolds are not generically area-minimizing, thus quantifying the failure of the Künneth formula at sending $\mathbb{Z}$-area-minimizing representatives to $\mathbb{Z}$-area-minimizing representatives.
Recall Assumptions 1.0.1 and 1.0.3. First of all, if $\Sigma]$ is a torsion class, then in many cases, $\Sigma]^{\prime}$ will annihilate $\Sigma]$ in their Cartesian product, i.e., $\Sigma]\times[\Sigma^{\prime}]=0$. To be precise, if $\kappa[\Sigma]=0$, then $\Sigma]\times[\Sigma^{\prime}]=0$ if $\Sigma^{\prime}]=\kappa[\Sigma^{\prime\prime}]$ for another class $\Sigma^{\prime\prime}]$. In these cases, it holds trivially that the Cartesian products fails to send $\mathbb{Z}$-area-minimizing representatives to $\mathbb{Z}$-area-minimizing representatives. The same reasoning holds for $\Sigma^{\prime}].$ Thus, we will only consider the case of both factors being free classes.
Theorem 4.
If $d\geq 1,c\geq 2$, $\Sigma]$ is a free class on an orientable $M$, then for any real number $\lambda>1,$ any metric $g^{\prime}$ on $M^{\prime}$, there exists a non-empty open set $\Omega_{[\Sigma],g^{\prime},\lambda}$ in the space of Riemannian metrics on $M$, a non-empty open set $\Omega_{[\Sigma],g^{\prime},\lambda}^{\prime}\ni g^{\prime}$ in the space of Riemannian metrics on $M^{\prime}$, and an integer $\nu_{[\Sigma],g^{\prime},\lambda}\geq 2$, such that for any metrics $h\in\Omega_{[\Sigma],g^{\prime},\lambda}$, $h^{\prime}\in\Omega_{[\Sigma],g^{\prime},\lambda}^{\prime}$ and any free homology class $\sigma]\in\nu H_{m}(M^{\prime},\mathbb{Z})$ with any integers $\nu\geq\nu_{[\Sigma],g^{\prime},\lambda},$ $0\leq m\leq d^{\prime}+c^{\prime},$ we have,
(1.0.14) $\left\lVert[\Sigma]\right\rVert_{h}^{\mathbb{Z}}\times\left\lVert[\sigma]\right\rVert_{h^{\prime}}^{\mathbb{Z}}>\lambda\left\lVert[\Sigma]\times[\sigma]\right\rVert_{h\times h^{\prime}}^{\mathbb{Z}}.$
We want to emphasize that in the above statement $\sigma]$ can be any free class of any dimension on $M^{\prime}$. It is an open problem [34] whether the above can hold among $R$-flat chains with $R=\mathbb{R}$ or $\mathbb{Z}/p\mathbb{Z}$ with $p$ a prime number.
The rest of this manuscript will be split into four parts. The first three provide the proof of Theorem 2, Theorem 3 and Theorem 4. Last, we will give some discussion of the theorems, prove some further results and give some further analogies with number theory.
Acknowledgements
The author would like thank his advisor Professor Camillo De Lellis for his unwavering support. Many helpful suggestions about this manuscript were due to him. Special tribute is paid to Professor William Allard and Professor Frederick Almgren. When the author was an undergraduate at Duke, Allard told many stories of the early days of geometric measure theory, including his conversation with Federer. One night the author had a dream, where Almgren was discussing flat chains with Allard, with Almgren emphasizing that “the key is to change the coefficients.” That dream sparked a long-standing quest to find the meaning of the enigmatic quote, a pursuit that ultimately culminated in this manuscript.
The author is deeply indebted to Professor Leon Simon for his extraordinary generosity, guidance, and kindness throughout the years. On the occasion of his 80th birthday, the author warmly extends best wishes to Professor Simon. Sincere thanks are paid to Professors Frank Morgan and Antoine Song, whose works [47, 32] are the inspirations for the author to go from local regularity to global geometry. Professor Morgan has also raised several suggestions that ultimately inspire Section 7.1. A conversation with Professor Michael Freedman partially inspired Theorem 3. The author would also like to thank Professor Chao Li, whose questions convince the author to write this manuscript. Special thanks go to Professor Yang Li, who have raised many helpful suggestions. He pointed out the example of Siegel mass formula and suggested the author to write the discussion around Theorem 4.
2. Basic Definitions and Technical Preliminaries
We will give some basic definitions and technical preliminaries in this section. The first four subsections are more or less standard definitions, and the experienced reader can skip them. The deeper technical preliminaries are in the last five subsections.
2.1. Basics of Homology
Recall Assumption 1.0.1. Fix a decomposition (1.0.1) of $H_{d}(M,\mathbb{Z})$ into subgroups, and construct a basis of $H_{d}(M,\mathbb{Z})$ with respect to this decomposition.
Assumption 2.1.1.
Let $\{f_{1},\cdots,f_{b_{d}},v_{1},\cdots,v_{|I|}\}$ be a basis of $H_{d}(M,\mathbb{Z})$ so that
  • $\{f_{1},\cdots,f_{b_{d}}\}$ is a $\mathbb{Z}$-linear independent basis of the factor $\mathbb{Z}^{b_{d}}$ in (1.0.1),
  • for all $j\in I,$ the element $v_{j}$ generates the $\mathbb{Z}/p_{j}^{\nu_{j}}\mathbb{Z}$ factor in (1.0.1).
We need the following lemma
Lemma 2.1.1.
If $X,Y\in H_{d}(M,\mathbb{Z})$ are two integral homology classes with the same $\mathbb{Z}/n\mathbb{Z}$-reduction, i.e., $X^{\mathbb{Z}/n\mathbb{Z}}=Y^{\mathbb{Z}/n\mathbb{Z}}$ and $\tau_{d}|n$, then XYX-Y is $n$ times a free class.
Proof.
It suffices to determine the kernel of $\mathbb{Z}/n\mathbb{Z}$-reduction. By [20, Section 3.2 p. 215], tensor products are distributive over direct sums and for any $m\in\mathbb{Z}_{\geq 2}$, we have
$\mathbb{Z}/m\mathbb{Z}\otimes\mathbb{Z}/n\mathbb{Z}=\mathbb{Z}/\operatorname{gcd}(m,n)\mathbb{Z}.$
By Definition 1.0.2 of $\tau_{d}$, we deduce that if $\tau_{d}|n,$ then $\operatorname{gcd}(p_{i}^{\nu_{i}},n)=p_{i}^{\nu_{i}}.$ Thus $\otimes\mathbb{Z}/n\mathbb{Z}$ is an isomorphism on the torsion subgroup of $H_{d}(M,\mathbb{Z}).$ This implies that the kernel of $\mathbb{Z}/n\mathbb{Z}$-reduction contains only free classes in $H_{d}(M,\mathbb{Z})$. Since the kernel of the map $\otimes\mathbb{Z}/n\mathbb{Z}:\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ is $n\mathbb{Z},$ by distributiveness of the tensor product, we are done.∎
We also need a classical result by Thom [48].
Lemma 2.1.2.
With the same assumptions as in Theorem 3, there exists an infinite subset $K\subset\mathbb{Z}_{\geq 2},$ such that for all $k\in K$, the integral homology class $k[\Sigma]$ can be represented by a smoothly embedded connected compact submanifold,
Proof.
First, since the codimension of $\Sigma]$ with respect to the ambient manifold $M$ is at least two, any smoothly embedded representative of $k[\Sigma]$ with finitely many connected components can be made into a connected submanifold by connected sum of submanifolds. The case of $d\geq 2$ can be found in [25, Lemma 2.8.2]. For $d=1$, this can be done by direct elementary geometry construction as follows. For two smoothly embedded circles $\gamma_{1},\gamma_{3}$, parametrized as smooth injections except at endpoints,
(2.1.1) $\gamma_{1}:[0,1]\to M,$
(2.1.2) $\gamma_{3}:[2,3]\to M,$
take a smooth curve
(2.1.3) $\gamma_{2}:[1,2]\to M,$
such that
$\gamma_{2}(1)=\gamma_{1}(1),\gamma_{2}(2)=\gamma_{3}(2).$
In other words, $\gamma_{2}$ goes from the endpoint of $\gamma_{1}$ to the start point of $\gamma_{2}.$ Define
$\gamma_{4}:[3,4]\to M,$
by
(2.1.4) $\gamma_{4}(t)=\gamma_{2}(1+(4-t)).$
Then $\gamma_{4}$ goes from the endpoint of $\gamma_{3}$ to the start point of $\gamma_{1}.$ Define $\gamma:[0,4]\to M$ by requiring the restriction of $\gamma$ to [i1,i][i-1,i] equals $\gamma_{i}$ for $i=1,2,3,4.$ Then $\gamma$ is a well-defined piecewise smooth closed curve. As integral currents, we have
$\gamma=\gamma_{1}+\gamma_{3},$
as by (2.1.4) $\gamma_{4}$ reversely traverses $\gamma_{2}$. Thus, $\gamma]=[\gamma_{1}]+[\gamma_{3}]$ as homology classes. Smoothing the corners of $\gamma$ and use transversality [50] by $c\geq 2$ to perturb $\gamma$ into a smoothly embedded curve $\gamma^{\prime}$, we can obtain a connected closed curve $\gamma^{\prime}$ representing the homology class $\gamma_{1}]+[\gamma_{2}].$
Thus, it suffices to prove the lemma without the connectedness assumption. Let us argue by contradiction. Suppose there exists only a finite subset $F\subset\mathbb{Z}_{\geq 2}$ so that for $k\in F$, the integral homology class has a smoothly embedded representative. By [48] some multiple of $\left(\prod_{k\in F}k\right)[\Sigma]$ admits a smoothly embedded representative, a contradiction. Here we regard $\prod_{k\in F}k$ as $1$ in case $F$ is empty. ∎
2.2. Flat Chain and Currents
Let us recall the discussion about flat chains before Definition 1.0.5. However, the space of flat chains is too large to have nice compactness properties. To this end, we need a smaller space. By [13, Sections 4.2.16 and (4.2.16)v of 4.2.26], we have,
Fact 2.2.1.
For $R=\mathbb{Z}$ or $\mathbb{Z}/n\mathbb{Z}$ for $n\in\mathbb{Z}_{\geq 2}$, finite mass $R$-flat chains form the space of $R$-rectifiable currents.
By [13, Sections 4.2.17 and (4.2.17)v of 4.2.26], the space of $R$-rectifiable currents with finite boundary mass enjoys compactness and closure properties that allow us to find $R$-area-minimizing representatives. Since we are concerned with area-minimization, it suffices to consider only $R$-rectifiable currents from now on for these two discrete $R.$
An alternative characterization for $\mathbb{Z}$-rectifiable currents [46, Section 27] is via the dual space of differential forms.
Fact 2.2.2.
Every $\mathbb{Z}$-rectifiable current $T$ belongs in the dual space of smooth differential forms, with actions on a form $\phi$ defined by
$T(\phi)=\int_{\operatorname{Supp}T}\theta_{T}\phi(\eta)d\mathcal{H}^{d},$
where $\operatorname{Supp}T$ is the Hausdorff $d$-dimensional (countably) underlying rectifiable set of $T$ defined in [13, Section 4.1.28(4)], $\theta_{T}$ is the density of the mass measure of $T$ and $\eta$ is the simple unit $d$-vector representing the oriented tangent space to $\operatorname{Supp}T$ Hausdorff $d$-dimensional a.e.
Conversely, we can also characterize $\mathbb{Z}$-flat chains using $\mathbb{Z}$ rectifiable currents. By [13, bottom of p.382 Section 4.1.24], we have,
Fact 2.2.3.
Every $d$-dimensional $\mathbb{Z}$-flat chain decomposes as a sum of a $d$-dimensional $\mathbb{Z}$-rectifiable current and the boundary of a $(d+1)$-dimensional $\mathbb{Z}$-rectifiable current and vice versa.
The situation with $\mathbb{R}$-flat chains is a bit more complicated. Analogues of Fact 2.2.1 and the corresponding compactness and closure theorems for $\mathbb{R}$-rectifiable currents no longer hold. However, by [14], the problem of finding $\mathbb{R}$-area-minimizing representatives can still be solved in $\mathbb{R}$-flat chains.
An important fact about $\mathbb{Z}/n\mathbb{Z}$ rectifiable current is as follows [13, p.430 of Section 4.2.26].
Fact 2.2.4.
For every $\mathbb{Z}/n\mathbb{Z}$-rectifiable current $T,$ there exists at least one $\mathbb{Z}$-rectifiable current $V$, called Federer’s ”representative” modulo $n$ of $T$ such that the $\mathbb{Z}/n\mathbb{Z}$-mass measure of $T$ coincides with the $\mathbb{Z}$-mass measure of $V$ and $T=V^{\mathbb{Z}/n\mathbb{Z}}.$
Unfortunately, Federer used the word ”representative” very differently from how we did. In order not to cause reader frustration, we will always use the phrase Federer’s ”representative” modulo $n$ together.
We want to emphasize that
Fact 2.2.5.
In general, Federer’s ”representative” modulo $n$
  • is not unique and many times there are infinitely many,
  • of a $\mathbb{Z}/n\mathbb{Z}$-cycle may not even be a $\mathbb{Z}$-cycle at all.
For instance, consider an embedded real projective plane $\mathbb{RP}^{2}$ inside the four-sphere $S^{4}.$ For each homotopically non-trivial embedded $C^{1}$ curve $\gamma$ on $\mathbb{RP}^{2}$, cut $\mathbb{RP}^{2}$ along $\gamma$ to obtain an orientable domain $\Delta.$ Then there exists a unique orientation of $\Delta$ so that $\partial\Delta=2\gamma$ as $\mathbb{Z}$-rectifiable currents. Any such $\Delta$ is a Federer’s ”representative” modulo $2$ of $\mathbb{RP}^{2}$, and thus illustrates both bullets of Fact 2.2.5.
2.3. Definition of regular and singular sets
Definition 2.3.1.
We say an $R$-rectifiable current $T$ is smooth at a point $p$ in the support of $T$ if there exists an open set $U$ containing $p$ on $M$, such that  $T$ restricted to $U$ equals $\alpha$ times a smooth submanifold $N$, with $\alpha\in R.$ The definitions of regular sets and singular sets of $T$ are as follows:
  • The singular set of $T,$ $\operatorname{Sing}T,$ is defined as set of points in the support of $T$ where $T$ is not smooth,
  • The regular set of $T,$ $\operatorname{Reg}T$, is defined as the set of points in the support of $T$ where $T$ is smooth.
Here, support means the underlying set of an integral current defined in [13, Section 4.1.28(4)]. From now on, we will also use the symbol $\operatorname{Supp}T$ to mean the support of $T.$ For example, the figure $8$ has a singular point at its self-intersection, while figure 0 counted with multiplicity 2-2 is smooth as $\mathbb{Z}/5\mathbb{Z}$-rectifiable current.
2.4. Almgren stratification of singular sets
Recall the definition of $d$-dimensional stationary varifolds in [1]. Roughly speaking, they can be understood as mass measures of unoriented countably rectifiable sets counted with positive integer multiplicities, and are critical points of $d$-dimensional area. As $\mathbb{Z}$-area-minimizing representatives and $\mathbb{Z}/n\mathbb{Z}$-area-minimizing representatives, with $n\in\mathbb{Z}_{\geq 2}$, are minimizers of $d$-dimensional area, their associated mass measures are naturally stationary varifolds.
For a $d$-dimensional stationary varifold $T$, the Almgren stratification of $T$ ([54]) is an ascending chain of closed subsets of the support of  $T$
$\mathcal{S}_{0}T\subset\mathcal{{S}}_{1}T\subset\cdots\subset\mathcal{S}_{d}T=\operatorname{Supp}T,$
such that $\mathcal{S}_{j}T$ consists of points in $\operatorname{Supp}T$ with tangent cones having at most $j$-dimensional translational invariance. The reader can just regard the Almgren stratification as a geometric measure theory analogue of Whitney stratification.
Define the $j$-th stratum of the Almgren stratification of $T$ as points in $\operatorname{Supp}T$ who has a tangent cone with precisely $j$-dimensional translation invariance. In other words,
Definition 2.4.1.
We call $\mathcal{S}_{j}T\setminus\mathcal{S}_{j-1}T$ the $j$-th stratum of the Almgren stratification, with $S_{-1}T=\emptyset.$
For integers $j\leq d-1,$ points in the $j$-th stratum of $T$ necessarily lie in the singular set of $T.$
The classical regularity result in [54] is
Fact 2.4.1.
The $j$-th Almgren stratification of $T$ has Hausdorff dimension at most $j.$
2.5. The norms across different metrics
In this subsection, we will collect several facts about the norms $\left\lVert\cdot\right\rVert^{R}_{g}$ across different metrics. First of all, let us enlarge the definition of $\left\lVert\cdot\right\rVert^{R}_{g}.$
Definition 2.5.1.
For any $R$-homology class $\omega\in H_{d}(M,R),$ define
(2.5.1) $\left\lVert\omega\right\rVert^{R}_{g}=\inf_{\begin{subarray}{c}T\in\mathcal{F}_{g}(M,R)\\ T\in\omega\end{subarray}}\mathbf{M}_{g}^{R}(T).$
Define the set of $R$-area-minimizing representatives $\mathbf{MIN}_{g}^{R}(\omega)$ of $\omega$ to be
$\mathbf{MIN}_{g}^{R}\left(\omega\right)=\left\{T|T\in\mathcal{F}_{g}(M,R),T\in\omega,\mathbf{M}^{R}_{g}(T)=\left\lVert\omega\right\rVert^{R}_{g}\right\}.$
Definition 1.0.7 and Definition 2.5.1  can be reconciled as
$\left\lVert[\Sigma]\right\rVert_{g}^{R}=$ $\left\lVert[\Sigma]^{R}\right\rVert^{R}_{g},$
$\mathbf{MIN}_{g}^{R}\left([\Sigma]\right)=$ $\mathbf{MIN}_{g}^{R}\left([\Sigma]^{R}\right).$
Using the argument in [46, 34.5 Theorem], we can show that
Fact 2.5.1.
$\left\lVert\cdot\right\rVert^{R}_{g}$ depends continuously on $g.$
For two different metrics $g,h$, by Definition 2.5.1, we have,
Fact 2.5.2.
If $\mathbf{M}_{g}^{R}\leq A\mathbf{M}_{h}^{R}$ for $A>0,$ then $\left\lVert\cdot\right\rVert_{g}^{R}\leq A\left\lVert\cdot\right\rVert_{h}^{R}$.
Using [14] and [13, Chapter 5], we have
Fact 2.5.3.
$\left\lVert\cdot\right\rVert^{R}_{g}$ is a norm in the sense of Definition 1.0.4, i.e., satisfying reversibility, triangle inequality and positivity.
A quick corollary of Fact 2.5.3 is that
Fact 2.5.4.
For any integer $k\in\mathbb{Z}$, $\omega\in H_{d}(M,R),$ we have
(2.5.2) $\left\lVert k\omega\right\rVert_{g}^{R}\leq|k|\left\lVert\omega\right\rVert_{g}^{R}.$
The inequality (2.5.2) can be strict in many situations, as we shall see in the proof of Corollary 1.0.2.
Now let us specialize to $R=\mathbb{R}$. Let us start with the basic facts [14]
Fact 2.5.5.
We have
  • $\left\lVert\cdot\right\rVert^{\mathbb{R}}_{g}$ is turns the the vector space $H_{d}(M,\mathbb{R})$ into a finite dimensional Banach space,
  • $\left\lVert[\Sigma]\right\rVert^{\mathbb{R}}_{g}\not=0$  if and only if $\Sigma]$ is a free class.
  • we have
    (2.5.3) $\lim_{\begin{subarray}{c}k\to\infty\\ k\in\mathbb{N}\end{subarray}}\frac{\left\lVert k[\Sigma]\right\rVert_{g}^{\mathbb{Z}}}{k}=\left\lVert[\Sigma]\right\rVert_{g}^{\mathbb{R}}.$
A specific corollary of the first bullet in Fact 2.5.5 is that $\left\lVert\cdot\right\rVert_{g}^{R}$ satisfies homogeneity, i.e., equality in (2.5.2) always holds for all  $k\in\mathbb{R}.$By the triangle inequality, a quick corollary of the second bullet is that
Fact 2.5.6.
For any torsion class $\tau\in H_{d}(M,\mathbb{Z})$, we have
$\left\lVert\tau+[\Sigma]\right\rVert^{\mathbb{R}}_{g}=\left\lVert[\Sigma]\right\rVert^{\mathbb{R}}_{g}.$
When $R=\mathbb{Z}/n\mathbb{Z}$, we have some partial control on the homogeneity of $\left\lVert\cdot\right\rVert_{g}^{\mathbb{Z}/n\mathbb{Z}}$.
Lemma 2.5.1.
If an integer $k\in\left(-\frac{n}{2},\frac{n}{2}\right]$ invertible in $\mathbb{Z}/n\mathbb{Z}$ with its unique inverse in $\left(-\frac{n}{2},\frac{n}{2}\right]$ denoted by $l,$ then for any $\omega\in H_{d}(M,\mathbb{Z}/n\mathbb{Z}),$ we have
$|l|^{-1}\left\lVert\omega\right\rVert_{g}^{\mathbb{Z}/n\mathbb{Z}}\leq\left\lVert k\omega\right\rVert_{g}^{\mathbb{Z}/n\mathbb{Z}}\leq|k|\left\lVert\omega\right\rVert_{g}^{\mathbb{Z}/n\mathbb{Z}},$
and thus
(2.5.4) $\frac{2}{n}\left\lVert\omega\right\rVert_{g}^{\mathbb{Z}/n\mathbb{Z}}\leq\left\lVert k\omega\right\rVert_{g}^{\mathbb{Z}/n\mathbb{Z}}\leq\frac{n}{2}\left\lVert\omega\right\rVert_{g}^{\mathbb{Z}/n\mathbb{Z}}.$
Proof.
Apply (2.5.2) to $k\omega$ and $l(k\omega)$. ∎
We need a basic lemma about general norms.
Lemma 2.5.2.
Fix a basis $\{v_{1},\cdots,v_{b}\}$ of $\mathbb{R}^{b}$. For any continuous Banach norm $\left\lVert\cdot\right\rVert$ on $\mathbb{R}^{n},$ there exists a $C>0$ that depends continuously on $\left\lVert\cdot\right\rVert,$ so that
$\sum_{j=1}^{b}|a_{j}|\left\lVert v_{j}\right\rVert\leq C\left\lVert\sum_{j=1}^{b}a_{j}v_{j}\right\rVert,$
with each $a_{j}\in\mathbb{R}.$
Proof.
For $v=\sum_{j=1}^{b}a_{j}v_{j}$ with each $a_{j}\in\mathbb{R}$ define
$\left\lVert v\right\rVert^{\infty}=\sum_{j=1}^{b}|a_{j}|\left\lVert v_{j}\right\rVert.$
It is straightforward to verify that $\left\lVert\cdot\right\rVert^{\infty}$ is a continuous function Banach norm on $\mathbb{R}^{n}$. We need to prove that $\frac{\left\lVert\cdot\right\rVert^{\infty}}$$\left\lVert\cdot\right\rVert$} is bounded. By homogeneity, it suffices to verify this on the norm unit sphere $\{v\in\mathbb{R}^{n}|\left\lVert v\right\rVert=1\}.$ This follows from continuity of $\left\lVert\cdot\right\rVert^{\infty},$ and the the compactness of norm unit sphere. ∎
Lemma 2.5.3.
There exists a continuous function $\upsilon$ mapping the space of Riemannian metrics to $\mathbb{R}_{\geq 0},$ such that
$\sup_{[\Sigma]\in H_{d}(M,\mathbb{Z})\textnormal{ is free}}\frac{\left\lVert[\Sigma]\right\rVert_{g}^{\mathbb{Z}}}{\left\lVert[\Sigma]\right\rVert_{g}^{\mathbb{R}}}\leq\upsilon(g).$
Proof.
Recall Assumption 2.1.1. Every free element $\Sigma]\in H_{d}(M,\mathbb{Z})$ can be written as
$[\Sigma]=\sum_{j=1}^{b_{d}}\alpha_{j}f_{j}+\sum_{i\in I}\beta_{i}v_{i},$
with each $\alpha_{j}\in\mathbb{Z},\beta_{i}\in\mathbb{Z}\cap[0,p_{j}^{\nu_{j}}],$ and not all $a_{i}$ are $0.$ By Lemma 2.5.2 there is a continuous function $C$ from the space of Riemannian metrics to $\mathbb{R}_{>0}$ so that
$\left\lVert[\Sigma]\right\rVert^{\mathbb{R}}_{g}=\left\lVert\sum_{j=1}^{b_{d}}\alpha_{j}f_{j}\right\rVert^{\mathbb{R}}_{g}\geq C^{-1}\sum_{j=1}^{b_{d}}|\alpha_{j}|\left\lVert f_{j}\right\rVert^{\mathbb{R}}_{g},$
where at the first equality we have used Fact 2.5.6. By the mediant inequality and $\max_{j}|\alpha_{j}|\geq 1$, we have
$\frac{\left\lVert[\Sigma]\right\rVert_{g}^{\mathbb{Z}}}{\left\lVert[\Sigma]\right\rVert_{g}^{\mathbb{R}}}\leq$ $C\frac{\sum_{j=1}^{b_{d}}|\alpha_{j}|\left\lVert f_{j}\right\rVert_{g}^{\mathbb{Z}}+\sum_{i\in I}|\beta_{i}|\left\lVert v_{i}\right\rVert^{\mathbb{Z}}_{g}}{\sum_{j=1}^{b_{d}}|\alpha_{j}|\left\lVert f_{j}\right\rVert^{\mathbb{R}}_{g}}$
$=$ $C\frac{\sum_{j=1}^{b_{d}}|\alpha_{j}|\left\lVert f_{j}\right\rVert_{g}^{\mathbb{Z}}}{\sum_{j=1}^{b_{d}}|\alpha_{j}|\left\lVert f_{j}\right\rVert^{\mathbb{R}}_{g}}+C\frac{\sum_{i\in I}|\beta_{i}|\left\lVert v_{i}\right\rVert^{\mathbb{Z}}_{g}}{\sum_{j=1}^{b_{d}}|\alpha_{j}|\left\lVert f_{j}\right\rVert^{\mathbb{R}}_{g}}$
$\leq$ $C\max_{1\leq j\leq b_{d}}\frac{\left\lVert f_{j}\right\rVert^{\mathbb{Z}}_{g}}{\left\lVert f_{j}\right\rVert_{g}^{\mathbb{R}}}+C\tau_{b}\frac{\max_{i\in I}\left\lVert v_{i}\right\rVert^{\mathbb{Z}}_{g}}{\min_{1\leq j\leq b_{d}}\left\lVert f_{j}\right\rVert_{g}^{\mathbb{R}}}.$
By Fact 2.5.1, we are done. ∎
2.6. Calibrations
The standard way to prove $\mathbb{Z}$-area-minimizing and $\mathbb{R}$-area-minimizing is via calibrations. The notion of calibrations started with the classical paper of Harvey-Lawson [18].
Recall that the comass of a  $d$-dimensional differential form  $\phi$ is the maximum of $\phi$ evaluated on simple unit $d$-vectors in the tangent space to $M$ among all points [13, Section 1.8]. In other words
(2.6.1) $\operatorname{comass}_{g}\phi=\max_{p\in M}\max_{\begin{subarray}{c}P\subset T_{p}M\\ \dim P=d\end{subarray}}\phi(P).$
Here we use the symbol $P$ to denote both a $d$-dimensional plane $P$ and the unit simple $d$-vector representing $P.$
For instance, if $\phi$ is a simple form, then the comass of $\phi$ is equal to the maximum Riemannian length of $\phi$ over all points.
Now we are ready to give the definition of calibrations.
Definition 2.6.1.
(Definition of calibrations)
  • We say a closed $d$-dimensional Hausdorff measurable $d$-form $\phi$ on a (possibly open) ambient manifold is a calibration if its comass is at most $1.$
  • We say a $d$-dimensional integral current $T$ is calibrated by $\phi,$ if $d$-dimensional Hausdorff measure almost everywhere $\phi$ restricted to the tangent space of $T$ equals the volume form of $T.$
In general, if the calibration form $\phi$ has sufficient regularity, then any integral current $T$ calibrated by $\phi$ will be area-minimizing. For smooth forms, Harvey-Lawson [18, Chapter II, Theorem 4.2] proved that
Fact 2.6.1.
If a $d$-dimensional integral current $T$ on a Riemannian manifold is calibrated by a degree $d$ smooth calibration form $\phi$, then $T$ is both $\mathbb{Z}$-area-minimizing and $\mathbb{R}$-area-minimizing, and any integral current homologous to $T$ and having the same area as $T$ must also be calibrated by $\phi.$
The widest possible notion of calibrations is found in $\mathbb{R}$-flat cochains [14], which are defined as bounded continuous linear functionals on $\mathbb{R}$-flat chains. By [15, Section 4.1.19], $\mathbb{R}$-flat cochains are equivalent to measurable forms with weak exterior derivatives that both have finite mass.
Definition 2.6.2.
We say an $\mathbb{R}$-flat cochain $\alpha,$ is a calibration on $(M,g),$ if
  • $d\alpha=0,$
  • $\alpha(T)\leq\mathbf{M}^{\mathbb{R}}_{g}(T),$ for all $\mathbb{R}$-flat chains $T.$
We say an $\mathbb{R}$-flat chain $T$ is calibrated by $\alpha,$ if
$\alpha(T)=\mathbf{M}^{\mathbb{R}}_{g}(T).$
By definition of calibrations, it is clear that any classical calibration form serves as a calibration $\mathbb{R}$-flat cochain via the canonical action of differential forms on $\mathbb{R}$-flat chains. Fact 2.6.1 holds when $T$ is an $\mathbb{R}$-flat chain and $\phi$ is a calibration $\mathbb{R}$-flat cochain.
By [14, Section 4.12], we have
Fact 2.6.2.
Every element of $\mathbf{MIN}_{g}^{\mathbb{R}}\left([\Sigma]\right)$ is calibrated by a calibration $\mathbb{R}$-flat cochain.
By Fact 2.6.1 and 2.6.2, we have
Fact 2.6.3.
A $\mathbb{Z}$-area-minimizing integral current representing $\Sigma]$ is calibrated by an $\mathbb{R}$-flat cochain if and only if $\left\lVert[\Sigma]\right\rVert_{g}^{\mathbb{Z}}=\left\lVert[\Sigma]\right\rVert_{g}^{\mathbb{R}}$.
2.7. Zhang’s constructions
We need a way to construct metrics so that we can make a smoothly embedded representative of an integral homology class area-minimizing.
Lemma 2.7.1.
Assume $N$ is a $d$-dimensional connected embedded compact closed submanifold of $M$. For any real number $\lambda>1,$  there exists a smooth tubular neighborhood $U(N)$ of $N,$ a smooth Riemannian metric $h$ on $M$, and a smooth deformation retract $\pi_{N}$ of $U(N)$ onto $N$, such that
  • any stationary varifold on $(M,h)$ whose support is not contained in $U(N)$ has area larger than $\lambda\mathbf{M}_{h}^{\mathbb{Z}}(N),$
  • $\pi_{N}^{\ast}\operatorname{dvol}_{N}$ is a calibration form on $(U(N),h),$
  • NN is the only integral current calibrated by $\pi_{N}^{\ast}\operatorname{dvol}_{N}$ in $(U(N),h)$,
  • if $N$ represents a non-zero $\mathbb{R}$-homology class, then $\pi_{N}^{\ast}\operatorname{dvol}_{N}$ can be extended to a smooth calibration form on $(M,h),$
  • NN is $\mathbb{Z}/n\mathbb{Z}$-area-minimizing in $U(N)$ for $n\in\mathbb{Z}_{\geq 2}$.
Proof.
The result is due to Yongsheng Zhang’s thesis [58] and published in [60, 59]. Take a tubular neighborhood $U(N)$ of $N$ inside $M.$ The first bullet is [59, Theorem 3.7]. The second to the fourth bullets are [60, Lemma 3.1] applied to $U(N)$ and $M$. The last bullet follows from [26, Lemma 2.12]. For an alternative proof the last bullet, by [25, Fact 7.0.6] $\pi^{\ast}_{N}\operatorname{dvol}_{N}$ is a calibration form on $U(N)$ implies that $\pi$ is a $d$-dimensional area-non-increasing map from $U(N)$ to $N$, i.e.,
(2.7.1) $\mathbf{M}^{R}(\pi_{\ast}T)\leq\mathbf{M}^{R}(T),$
for $R=\mathbb{Z},$ or $\mathbb{Z}/n\mathbb{Z}$ with $n\in\mathbb{Z}_{\geq 2}$ and $T$ any $R$-flat chains supported in $U(N).$ For any $R$-flat chain $N^{\prime}$ homologous to $N$ in $U(N)$, we necessarily have $\pi_{\ast}N^{\prime}=N$ since $\pi$ is a deformation retract. By (2.7.1), we are done. ∎
2.8. Monotonicity formula across metrics
In this subsection, we prove a monotonicity formula for stationary varifolds that holds across different metrics. Recall Assumption 1.0.1.
Lemma 2.8.1.
There exist an open set $\Omega_{g}$ containing $g$ in the space of smooth Riemannian metrics, positive real numbers $C_{g},c_{g},r_{g}>0$ depending on $g$,  such that, for any metric $h\in\Omega_{g}$, any point $p\in M$ and  all radii $r\in(0,r_{g}],$ we have
  • for any $d$-dimensional integral stationary varifold $V$ on $M$ in the metric $h$, the function,
    $\exp(c_{g}r)V(B_{r}(p,h))r^{-d},$
    is monotonically increasing in $r$, where $V(\cdot)$ denotes the mass measure of $V$, and $B_{r}(p,h)$ is the radius $r$ geodesic ball in metric $h$  centered at $p,$
  • and the density $\theta_{V}({p,h})$ of $V$ at $p$ in $h$ is well-defined and satisfies
    $\theta_{V}({p,h})<C_{g}V(M).$
Proof.
This is folklore and we will only give a sketch of the proof. The argument is the Riemannian adaptation of the first two sections of [5]. First by [11, Section 8 Theorem], the injectivity radius $\operatorname{Inj}_{g}(M)$ of $M$ in $g$ depends continuously on the metric $g$. Thus, there exists an open neighborhood $\Omega_{g}$ containing $g$ in the space of Riemannian metrics so that
$r_{g}=\frac{1}{2}\inf_{h\in\Omega_{g}}\operatorname{Inj}_{h}(M)>0.$
For any metric $h\in\Omega_{g},$ and any point $p\in M$, adopt a normal coordinate $(x_{1},\cdots,x_{d+c})$ in $B_{r_{g}}(p,h)$. Let $\mathbf{r}_{p,h}$ denote the distance function to $p$ on $(M,h)$. By [17, Chapter 2 ] we have
(2.8.1) $\mathbf{r}_{p,h}\nabla\mathbf{r}_{p,h}=\sum_{j}x_{j}\partial_{j}.$
A straightforward calculation using Taylor expansion shows that for any unit length tangent vector $u$ in $B_{r_{g}}(p,h)$, we have
(2.8.2) $|\langle\nabla_{u}(\mathbf{r}_{p,h}\nabla\mathbf{r}_{p,h}),u\rangle-1|\leq c_{p,h},$
for some $c_{p,h}>0$ which depends only on the $C^{3}$ norm of the ambient metric $h$, i.e.,
(2.8.3) $c_{p,h}=O(\left\lVert h\right\rVert_{C^{3}}).$
By compactness of $M$ and shrinking $\Omega_{g}$ if necessary, we can set
$c_{g}=\sup_{p\in M,h\in\Omega_{g}}c_{p,h}<\infty.$
Arguing as in the first two sections of [5], we deduce from (2.8.2) that $\exp(c_{g}r)V(B_{r}(p,h))r^{-d}$ is monotonically increasing in $r$.
The second bullet follows from the first bullet and
(2.8.4) $\theta_{V}({p,h})=\lim_{r\to 0}\frac{V(B_{r}(p,h))}{r^{d}}\leq\exp(c_{g}r_{g})V(B_{r_{g}}(p,h))r_{g}^{-d}\leq C_{g}V(M).$
As a direct corollary, we deduce that
Lemma 2.8.2.
Using the same notation as Lemma 2.8.1, for all $n\geq 2,p\in M,h\in\Omega_{g}$ and any $T\in\mathbf{MIN}_{h}^{\mathbb{Z}/n\mathbb{Z}}\left([\Sigma]\right),$ we have
$\theta_{T}(p,h)\leq C_{[\Sigma]},$
with $C_{[\Sigma]}>0$ depending only on $\Sigma]$ and $\Omega_{g}.$
Proof.
Use Lemma 2.8.1 and Fact 1.0.2. ∎
2.9. Regularity theorems
In this subsection, we will prove an essential lemma about the regularity of $\mathbb{Z}/n\mathbb{Z}$ area-minimizing representatives. Recall Fact 2.2.4.
Lemma 2.9.1.
Suppose $T$ is a $\mathbb{Z}/n\mathbb{Z}$-rectifiable current representing a $\mathbb{Z}/n\mathbb{Z}$-homology class. Let $S$ be a rectifiable current that is a Federer’s ”representative” modulo $n$ of $T$. Then
$\operatorname{Supp}\partial S\subset\left(\operatorname{Reg}(T)\cap\left\{p\bigg|\theta_{T}(p)<\frac{n}{2}\right\}\right)^{\complement},$
where $\operatorname{Reg}(T)$ denotes the regular part of $T.$
Here $\theta_{T}(p)$ is the density of $T$ at $p.$
Proof.
We will prove that if $p\in\operatorname{Supp}\partial S\cap\textnormal{Reg}(T),$ then $\theta_{T}(p)=\frac{n}{2}.$
Since $p$ is in the regular set of $T,$ there exists a neighborhood $B_{r}(p)$ in which $T$ restricted to $B_{r}(p)$ equals $\theta_{T}(p)N$ for some smooth submanifold $N$ of $B_{r}(p)$, so that in some coordinate system $(x_{1},\cdots,x_{d+c})$, $N$ is a smooth ball contained in the $x_{1}\cdots x_{d}$-plane. By [57, Corollary 1.4], there exists a $\mathbb{Z}$-flat chain $R$ in $B_{r}(p)$ so that
[…]
If $R=0,$ then $\partial S$ restricted to $B_{r}(p)$ is empty, as $S$ coincides with $\theta_{T}(p)N.$ Thus, we only have to deal with the case of $R\not=0.$
Recall Fact 2.2.2. There exist a $d$-dimensional rectifiable set $B$, an $\mathcal{H}^{d}$ measurable simple unit $d$-vector field $\eta$ , so that the action of the current $R$ on differential forms is
$R(\phi)=\int_{B}\theta_{R}\phi(\eta)d\mathcal{H}^{d}(x),$
with $\phi$ any smooth $d$-dimensional form.
By [13, page 430], the support of $S$ restricted to $B_{r}(p)$ is contained in $N,$ thus by (2.9.1) the support of $R$ is also contained in $N,$ i.e., $B\subset N.$ This implies that $\eta=\pm\mathbf{T}_{x}N$ Hausdorff $d$-dimensional a.e. Here $\mathbf{T}_{x}N$ means the unique simple $d$-vector field representing the tangent space at $x$ to $N.$
Thus, we have
[…]
By [13, Section 4.1.28], we deduce that
[…]
However, since we have $R\not=0$, i.e., $\theta_{R}(x)$ being a positive integer on $B$ Hausdorff $d$-dimensional a.e., and $\theta_{T}(p)\leq\frac{n}{2}$  by [13, page 430], we have
Fact 2.9.1.
The inequality
(2.9.2) $|\theta_{T}(p)\pm n\theta_{R}(x)|\geq\frac{n}{2},$
holds where equality is achieved if and only if $\theta_{T}(p)=\frac{n}{2},\theta_{R}(x)=1.$
Thus, we have
[…]
By Fact 2.2.4, we  have
[…]
In other words, equality holds in (2.9.2) Hausdorff $d$-dimensional a.e. on $B.$ By Fact 2.9.1, we deduce that
$\theta_{T}(p)=\frac{n}{2},\theta_{R}(x)=1,$
on $B$ Hausdorff $d$-dimensional a.e. Thus $\partial S$ intersect $\operatorname{Reg}T$ at $\theta_{T}=\frac{n}{2}$ and we are done. ∎
3. Proof of Theorem 2 and Corollary 1.0.1
Now we are ready to prove Theorem 2 and Corollary 1.0.1.
3.1. Any Federer’s ”representative” modulo $n$ of an element in $\mathbf{MIN}_{g}^{\mathbb{Z}/n\mathbb{Z}}\left([\Sigma]\right)$ is an integral cycle for $n$ large
Recall Lemma 2.8.2. For ambient metric $h\in\Omega_{g}$, let $T\in\mathbf{MIN}_{h}^{\mathbb{Z}/n\mathbb{Z}}\left([\Sigma]\right)$ with $n\geq 3C_{[\Sigma]}$. Then for any point $p\in M,$ we have
(3.1.1) $\theta_{T}(p,h)<\frac{n}{2}.$
For any $\mathbb{Z}$-rectifiable current $S$ which is a Federer’s ”representative” modulo $n$   of $T$, by Lemma 2.9.1, we have
(3.1.2) $\operatorname{Supp}\partial S\subset\operatorname{Sing}T.$
We claim that
Lemma 3.1.1.
$\operatorname{Sing}T$ has Hausdorff dimension at most (d2).(d-2).
Proof.
Recall Section 2.4. Let us consider stratum by stratum in the Almgren stratification of $T.$ By [10, Theorem 1.6], the $d$-th stratum of the Almgren stratification of $\operatorname{Sing}T$ is empty. By [13, Section 5.4.8], at any point $p$ in the (d1)(d-1)-stratum of the Almgren stratification of $\operatorname{Sing}T,$ the tangent cone $C_{p}$ to $T$ at $p$ splits as the product of $\mathbb{R}^{d-1}$ with a $1$-dimensional $\mathbb{Z}/n\mathbb{Z}$-length minimizing cone. By elementary geometry, it is straightforward to verify that any $1$-dimensional $\mathbb{Z}/n\mathbb{Z}$-length minimizing cone is a union of $kn$ rays (counted with multiplicity) starting from the origin for $k\in\mathbb{Z}_{\geq 1}$, which implies that $\theta_{T}(p,h)\geq\frac{n}{2}$. This is impossible, so the (d1)(d-1)-th stratum of the Almgren stratification is also empty. By Fact 2.4.1, we are done. ∎
By [13, Section 4.1.20], a (d1)(d-1)-dimensional $\mathbb{Z}$-flat chain supported in a (d1)(d-1)-dimensional integral geometric measure zero set must equal zero. Since Hausdorff measure zero implies integral geometric measure zero by [13, Section 2.10.6], a (d1)(d-1)-dimensional $\mathbb{Z}$-flat chain supported in a (d1)(d-1)-dimensional Hausdorff measure zero set must equal to zero. By Fact 2.2.3 and Lemma 3.1.1 applied to (3.1.2), we deduce that $\partial S$ is a (d1)(d-1)-dimensional $\mathbb{Z}$-flat chain supported on a set of Hausdorff dimension at most (d2)(d-2). Thus, we deduce that $\partial S=0.$
In other words, we have proved that
Fact 3.1.1.
Any $\mathbb{Z}$-rectifiable current $S$ that is Federer’s ”representative” modulo $n$ of $T$ is a $\mathbb{Z}$-cycle provided $n$ is large enough.
The above fact is precisely Corollary 1.0.1.
3.2. Determine the integral homology class of $S$
Since $S$ is a cycle, it is an integral current and belongs to a homology class $\Sigma^{\prime}].$ Let us first note that
Fact 3.2.1.
$S$ is $\mathbb{Z}$-area-minimizing in $\Sigma^{\prime}].$
Proof.
Let $W$ be another integral current in $\Sigma^{\prime}].$ By definition of Federer’s ”representative” modulo $n$ (Fact 2.2.4), we have
(3.2.1) $\mathbf{M}_{h}^{\mathbb{Z}}(S)=\mathbf{M}_{h}^{\mathbb{Z}/n\mathbb{Z}}(T).$
On the other hand, by area-minimality of $T$ and that the canonical homomoprhism $\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ is norm non-increasing, we deduce that
(3.2.2) $\mathbf{M}_{h}^{\mathbb{Z}/n\mathbb{Z}}(T)\leq\mathbf{M}_{h}^{\mathbb{Z}/n\mathbb{Z}}(W^{\mathbb{Z}/n\mathbb{Z}})\leq\mathbf{M}_{h}^{\mathbb{Z}}(W).$
Combining (3.2.1) and (3.2.2), we are done. ∎
By Fact 3.2.1 and (3.2.1), we deduce that
(3.2.3) $\left\lVert[\Sigma]\right\rVert^{\mathbb{Z}/n\mathbb{Z}}_{h}=\left\lVert[\Sigma^{\prime}]\right\rVert_{h}^{\mathbb{Z}}.$
Expand $\Sigma^{\prime}],[\Sigma]$ with respect to the basis in Assumption 2.1.1, we have
(3.2.4) $[\Sigma^{\prime}]=\sum_{j=1}^{b_{d}}\gamma_{j}f_{j}+\sum_{i\in I}\eta_{i}v_{i},$
and
(3.2.5) $[\Sigma]=\sum_{j=1}^{b_{d}}\alpha_{j}f_{j}+\sum_{i\in I}\beta_{i}v_{i},$
with each $\alpha_{j},\gamma_{j}\in\mathbb{Z},\beta_{i},\eta_{i}\in\mathbb{Z}\cap[0,p_{j}^{\nu_{j}}].$
Now assume $\tau_{d}|n$, which is assumed in Theorem 2. From $\Sigma]^{\mathbb{Z}/n\mathbb{Z}}=[\Sigma^{\prime}]^{\mathbb{Z}/n\mathbb{Z}},$ and Lemma 2.1.1, we deduce that for $1\leq j\leq b_{d},i\in I$ we have
$\gamma_{j}=$ $\alpha_{j}+k_{j}n,$
$\eta_{i}=$ $\beta_{i},$
with $k_{i}\in\mathbb{Z}.$ Thus, by Fact 1.0.2, Fact 2.5.6, and  Lemma 2.5.2, we have
$\left\lVert[\Sigma^{\prime}]\right\rVert_{h}^{\mathbb{Z}}\geq$ $\left\lVert[\Sigma^{\prime}]\right\rVert_{h}^{\mathbb{R}}$
$=$ $\left\lVert\sum_{j=1}^{b_{d}}\gamma_{j}f_{j}\right\rVert^{\mathbb{R}}_{h}$
$=$ $\left\lVert\sum_{j=1}^{b_{d}}(\alpha_{j}+k_{j}n)f_{j}\right\rVert^{\mathbb{R}}_{h}$
$\geq$ $(C(h))^{-1}\sum_{j=1}^{b_{d}}\left\lVert(\alpha_{j}+k_{j}n)f_{j}\right\rVert^{\mathbb{R}}_{h}$
$\geq$ $(C(h))^{-1}\sum_{j=1}^{b_{d}}\big(|k_{j}|n-|\alpha_{j}|\big)\left\lVert f_{j}\right\rVert^{\mathbb{R}}_{h}.$
where $C(h)$ is a continuous function from the space of Riemannian metrics to $\mathbb{R}_{>0}$. By (3.2.3) and (1.0.2), we deduce that
(3.2.6) $\left\lVert[\Sigma]\right\rVert_{h}^{\mathbb{Z}}\geq(C(h))^{-1}\sum_{j=1}^{b_{d}}\big(|k_{j}|n-|\alpha_{j}|\big)\left\lVert f_{j}\right\rVert^{\mathbb{R}}_{h}.$
By shrinking $\Omega_{g}$ to a smaller open set $\Omega_{[\Sigma],g}$, we can assume that
$\inf_{h\in\Omega_{g}}C(h)^{-1}>$ $0,$
$\inf_{h\in\Omega_{g}}\min_{1\leq j\leq b_{d}}\left\lVert f_{j}\right\rVert^{\mathbb{R}}_{h}>$ $0,$
$\sup_{h\in\Omega_{g}}\left\lVert[\Sigma]\right\rVert^{\mathbb{Z}}_{h}<$ $\infty.$
Definition 3.2.1.
Set
$N_{[\Sigma],g}=3\max\left\{C_{[\Sigma]},\frac{\sup_{h\in\Omega_{g}}\left\lVert[\Sigma]\right\rVert^{\mathbb{Z}}_{h}}{\left(\inf_{h\in\Omega_{g}}(C(h))^{-1}\right)\left(\inf_{h\in\Omega_{g}}\min_{1\leq j\leq b_{d}}\left\lVert f_{j}\right\rVert^{\mathbb{R}}_{h}\right)}+\max_{1\leq j\leq b_{d}}|\alpha_{j}|\right\}.$
Apply Definition 3.2.1 for $n\geq N_{[\Sigma],g}$. If $k_{j}\not=0,$ we have
$(C(h))^{-1}\sum_{j=1}^{b_{d}}\big(|k_{j}|n-|\alpha_{j}|\big)\left\lVert f_{j}\right\rVert^{\mathbb{R}}_{h}$
$\geq$ $3\frac{(C(h))^{-1}\left\lVert f_{j}\right\rVert^{\mathbb{R}}_{h}}{\left(\inf_{h\in\Omega_{g}}(C(h))^{-1}\right)\left(\inf_{h\in\Omega_{g}}\min_{1\leq l\leq b_{d}}\left\lVert f_{l}\right\rVert^{\mathbb{R}}_{h}\right)}\sup_{h\in\Omega_{g}}\left\lVert[\Sigma]\right\rVert^{\mathbb{Z}}_{h}$
$+2(C(h))^{-1}\left\lVert f_{j}\right\rVert^{\mathbb{R}}_{h}\max_{1\leq l\leq b_{d}}|\alpha_{l}|$
$\geq$ $3\left\lVert[\Sigma]\right\rVert^{\mathbb{Z}}_{h}.$
By (3.2.6), we deduce that $\sum_{j=1}^{b_{d}}|k_{j}|=0.$ In other words, we have
Fact 3.2.2.
For $n\geq N_{[\Sigma],g},$ with $\tau_{d}|n,$ we have $\Sigma]=[\Sigma^{\prime}]$.
3.3. Proof of Theorem 2
The first bullet of Theorem 2 follows from Fact 3.2.2. To prove the second bullet, note that everything follows if we can show that $S$ as a Federer’s ”representative” modulo $n$ is unique. Let $S^{\prime}$ be another Federer’s ”representative” modulo $n$ of $T.$
Lemma 3.3.1.
We have $S=S^{\prime}$.
Proof.
Recall Fact 3.2.2, $S$ and $S^{\prime}$ are homologous and have the same support.
By Lemma 3.1.1 and [13, Section 4.1.20], we know that the regular set of $S,S^{\prime}$ consists of countably (possibly open) pairwise disjoint boundaryless oriented submanifolds $X_{0},X_{1},X_{2},\cdots,$ such that
$\operatorname{Supp}S=\overline{\cup_{j\in\mathbb{N}}X_{j}}.$
By the constancy theorem [36, Section 4.9], we deduce that $S,S^{\prime}$ can be written as the chain sum
$S=$ $\sum_{j\in\mathbb{N}}a_{j}X_{j},$
$S^{\prime}=$ $\sum_{j\in\mathbb{N}}b_{j}X_{j},$
where each $a_{j},b_{j}\in\mathbb{N}.$
From $(S-S^{\prime})^{\mathbb{Z}/v\mathbb{Z}}=0,$ we deduce that
$\sum_{j\in\mathbb{N}}(a_{j}-b_{j})X_{j}^{\mathbb{Z}/n\mathbb{Z}}=0.$
However, as the collection of $X_{j}$ is pairwise disjoint, we must have $(a_{j}-b_{j})X_{j}^{\mathbb{Z}/n\mathbb{Z}}=0$ for each $j\in\mathbb{N},$ which by definition of $\left\lVert\cdot\right\rVert^{\mathbb{Z}/n\mathbb{Z}}$ only happens when
$a_{j}\equiv b_{j}(\mod n).$
By (3.1.1), this implies that $a_{j}=b_{j}$ for all $j\in\mathbb{N}.$ We are done. ∎
4. Proof of Theorem 3 and Corollary 1.0.2
In this section, we will prove Theorem 3 and Corollary 1.0.2. Let us start with Corollary 1.0.2, which is easier.
4.1. Proof of Corollary 1.0.2
Let us start with the following fact
Fact 4.1.1.
A $\mathbb{Z}$-area-minimizing current in a torsion homology class can never be calibrated.
This follows from the definition of $\mathbb{R}$-homology, as any closed $\mathbb{R}$-flat cochain $\alpha$ must evaluate to 0 on a torsion class.
Thus, to prove Corollary 1.0.2, we only need to consider free classes. Corollary 1.0.2 follows directly from the following lemma.
Lemma 4.1.1.
If $d\geq 1,c\geq 2,$ $\Sigma]$ is a free class and $M$ is an orientable, then for any $\lambda>2$, there exists a non-empty open set $\Omega^{\prime}_{[\Sigma],\lambda}$ in the space of Riemannian metrics, depending on $\Sigma]$ and $\lambda,$ such that for any metric $h\in\Omega^{\prime}_{[\Sigma],\lambda},$ we have
(4.1.1) $\left\lVert[\Sigma]\right\rVert_{h}^{\mathbb{Z}}>\lambda\mu_{[\Sigma]}\left\lVert[\Sigma]\right\rVert^{\mathbb{R}}_{h}.$
We emphasize that by Definition 1.0.1, we have $\mu_{[\Sigma]}\geq 2.$
Proof.
By Fact 2.5.1, it suffices to show the existence of one metric $h$ in which inequality (4.1.1) holds.
Let $N$ be a smoothly embedded connected submanifold representing $\mu_{[\Sigma]}[\Sigma].$ Apply Lemma 2.7.1 to $N$ to obtain a metric $h$ on $M,$ tubular neighborhood $U(N)$, and the deformation retract $\pi_{N}.$
By the last bullet of Lemma 2.7.1 and Fact 2.6.1, we have
(4.1.2) $\left\lVert\mu_{[\Sigma]}[\Sigma]\right\rVert_{h}^{\mathbb{Z}}=\left\lVert\mu_{[\Sigma]}[\Sigma]\right\rVert_{h}^{\mathbb{R}}=\mathbf{M}_{h}^{\mathbb{Z}}(N).$
By the first bullet of Fact 2.5.5, we have
(4.1.3) $\left\lVert[\Sigma]\right\rVert_{h}^{\mathbb{R}}=\frac{1}{\mu_{[\Sigma]}}\mathbf{M}_{h}^{\mathbb{Z}}(N).$
Let $T\in\mathbf{MIN}_{h}^{\mathbb{Z}}\left([\Sigma]\right).$ Since $\pi$ is a deformation retract of $U(N)$ onto $N,$ by homotopy invariance of homology [20], we deduce that $H_{d}(U(N),\mathbb{Z})=\mathbb{Z}[N]=\mathbb{Z}[\mu_{[\Sigma]}\Sigma].$ Thus, any integral current whose support is contained in $U(N)$ is homologous to some integer multiple of $\mu_{[\Sigma]}[\Sigma].$ Since $\Sigma]$ is a free class, we deduce that
Fact 4.1.2.
The support of $T$ cannot be contained in $U(N)$.
By the first bullet of Lemma 2.7.1, we deduce that
(4.1.4) $\left\lVert[\Sigma]\right\rVert_{h}^{\mathbb{Z}}>\lambda\mathbf{M}_{h}^{\mathbb{Z}}(N).$
Combining (4.1.3) and (4.1.4). We are done ∎
4.2. Proof of Theorem 3
The proof of Theorem 3 is reliant on the proof of Lemma 4.1.1.
By the last bullet of Lemma 2.7.1, NN is $\mathbb{Z}/n\mathbb{Z}$-area-minimizing in $U(N)$. By the first bullet of Lemma 2.7.1, we deduce that
$\left\lVert\mu_{[\Sigma]}[\Sigma]\right\rVert_{g}^{\mathbb{Z}/n\mathbb{Z}}=\mathbf{M}_{h}^{\mathbb{Z}}(N).$
If $n\geq 2$ and $\operatorname{gcd}\left(n,\mu_{[\Sigma]}\right)=1,$ then $\mu_{[\Sigma]}$ is invertible in $\mathbb{Z}/n\mathbb{Z}.$ By Lemma 2.5.1, we have
(4.2.1) $\frac{2}{n}\left\lVert[\Sigma]\right\rVert_{g}^{\mathbb{Z}/n\mathbb{Z}}\leq\left\lVert\mu_{[\Sigma]}[\Sigma]\right\rVert_{g}^{\mathbb{Z}/n\mathbb{Z}}=\mathbf{M}_{h}^{\mathbb{Z}}(N).$
Combining (4.2.1) and (4.1.4), we have
$\left\lVert[\Sigma]\right\rVert_{h}^{\mathbb{Z}}>\frac{2\lambda}{n}\left\lVert[\Sigma]\right\rVert_{h}^{\mathbb{Z}/n\mathbb{Z}}.$
By Fact 2.5.1, we are done by setting $\rho=\frac{2\lambda}{n}$ and $\Omega^{\prime}_{[\Sigma],n,\rho}=\Omega^{\prime}_{[\Sigma],\lambda}$ using the same notations as in Lemma 4.1.1.
5. Proof of Theorem 4
Let us first prove that,
Lemma 5.0.1.
For any metric $h$ on $M,$ there exists a non-empty open set $\Omega_{[\Sigma],h}\ni h$ in the space of Riemannian metrics on $M$, and an integer $\nu_{[\Sigma],h}$ such that for any integer $\nu\geq\nu_{[\Sigma],h}$ and any metric ${\hat{h}}\in\Omega_{[\Sigma],h}$, we have
(5.0.1) $1\leq\frac{\left\lVert\nu[\Sigma]\right\rVert^{\mathbb{Z}}_{{\hat{h}}}}{\left\lVert\nu[\Sigma]\right\rVert^{\mathbb{R}}_{{\hat{h}}}}\leq\frac{3}{2}.$
Proof.
By (2.5.3), there exists an integer $\nu_{[\Sigma],h}$ such that for any $\nu\geq\nu_{[\Sigma],h}$, we have
(5.0.2) $1\leq\frac{\left\lVert\nu[\Sigma]\right\rVert^{\mathbb{Z}}_{h}}{\left\lVert\nu[\Sigma]\right\rVert^{\mathbb{R}}_{h}}\leq\frac{9}{8}.$
There also exists an open set $\Omega_{[\Sigma],h}\ni h$ in the space of Riemannian metrics on $M$, such that for any metric ${\hat{h}}\geq\Omega_{[\Sigma],h}$, we have
(5.0.3) $\left(\frac{8}{9}\right)^{\frac{2}{d}}{\hat{h}}\leq h\leq\left(\frac{9}{8}\right)^{\frac{2}{d}}{\hat{h}}$,
as quadratic forms, which by definition of mass [13, p.358], implies that
(5.0.4) $\frac{8}{9}\mathbf{M}_{{\hat{h}}}^{R}\leq\mathbf{M}_{h}^{R}\leq\frac{9}{8}\mathbf{M}_{{\hat{h}}}^{R},$
for $R=\mathbb{Z}$ or $\mathbb{R}$. Combining (5.0.4) and (5.0.2), by Fact 2.5.2, we have
$\frac{\left\lVert\nu[\Sigma]\right\rVert^{\mathbb{Z}}_{{\hat{h}}}}{\left\lVert\nu[\Sigma]\right\rVert^{\mathbb{R}}_{{\hat{h}}}}\leq\frac{\frac{9}{8}\left\lVert\nu[\Sigma]\right\rVert^{\mathbb{Z}}_{h}}{\frac{8}{9}\left\lVert\nu[\Sigma]\right\rVert^{\mathbb{R}}_{h}}\leq\left(\frac{9}{8}\right)^{3}.$
Since $\left(\frac{9}{8}\right)^{3}<\frac{3}{2}$, we are done. ∎
Let us first find the non-empty open set of metrics on $M^{\prime}$. Apply Lemma 2.5.3. There exists a non-empty open set $\Omega_{[\Sigma],g^{\prime},\lambda}^{\prime}\ni g^{\prime}$ in the space of Riemannian metrics on $M^{\prime}$, such that for any $h^{\prime}\in\Omega_{[\Sigma],g^{\prime},\lambda}^{\prime}$, and any free class $\sigma]$ of any dimension $0\leq m\leq d^{\prime}+c^{\prime},$ we have,
(5.0.5) $\left\lVert[\sigma]\right\rVert_{h^{\prime}}^{\mathbb{R}}\leq\left\lVert[\sigma]\right\rVert_{h^{\prime}}^{\mathbb{Z}}\leq v_{g^{\prime}}\left\lVert[\sigma]\right\rVert_{h^{\prime}}^{\mathbb{R}},$
for some real number $v_{g^{\prime}}>0.$
Next let us find the non-empty open set of metrics on $M$. Apply Lemma 4.1.1. For any $\lambda>1$, there exists a non-empty open set $\Omega_{[\Sigma],\lambda}$ in the space of Riemannian metrics on $M$ such that for any metric $\check{h}$ in $\Omega_{[\Sigma],\lambda}$, we have $\left\lVert[\Sigma]\right\rVert_{\check{h}}^{\mathbb{Z}}>\lambda\mu_{[\Sigma]}\left\lVert[\Sigma]\right\rVert^{\mathbb{R}}_{\check{h}}.$
Fix one metric $\check{h}$ in $\Omega_{[\Sigma],\lambda}$. Apply Lemma 5.0.1 to obtain a non-empty open subset $\Omega_{[\Sigma],\lambda,\check{h}}\ni\check{h}$ contained in $\Omega_{[\Sigma],\lambda}$, and an integer $\nu_{[\Sigma],\check{h}}$, such that for any metric $h\in\Omega_{[\Sigma],\lambda,\check{h}}$, and any integer $\nu\geq\nu_{[\Sigma],\check{h}}$, we always have,
(5.0.6) $\left\lVert\nu[\Sigma]\right\rVert_{h}^{\mathbb{R}}\leq\left\lVert\nu[\Sigma]\right\rVert^{\mathbb{Z}}_{h}\leq\frac{3}{2}\left\lVert\nu[\Sigma]\right\rVert_{h}^{\mathbb{R}},$
and
(5.0.7) $\left\lVert[\Sigma]\right\rVert^{\mathbb{Z}}_{h}>\lambda\mu_{[\Sigma]}\left\lVert[\Sigma]\right\rVert_{h}^{\mathbb{R}}.$
For any integer $\nu\geq\nu_{[\Sigma],\check{h}}$, if $\sigma]=\nu[\sigma^{\prime}]\in\nu H_{m}(M,\mathbb{Z})$ is a free class. In these cases, using (1.0.12), (5.0.5), and (5.0.6), we have
(5.0.8) $\left\lVert[\Sigma]\times[\sigma]\right\rVert^{\mathbb{Z}}_{h}=\left\lVert[\Sigma]\times\nu[\sigma^{\prime}]\right\rVert_{h}^{\mathbb{Z}}$
(5.0.9) $=$ $\left\lVert\nu[\Sigma]\times[\sigma^{\prime}]\right\rVert_{h}^{\mathbb{Z}}$
(5.0.10) $\leq$ $\left\lVert\nu[\Sigma]\right\rVert_{h}^{\mathbb{Z}}\times\left\lVert[\sigma^{\prime}]\right\rVert_{h}^{\mathbb{Z}}$
(5.0.11) $\leq$ $\frac{3}{2}\left\lVert\nu[\Sigma]\right\rVert_{h}^{\mathbb{R}}\times v_{g^{\prime}}\left\lVert[\sigma^{\prime}]\right\rVert_{h}^{\mathbb{R}}$
(5.0.12) $=$ $\frac{3}{2}v_{g^{\prime}}\nu\left\lVert[\Sigma]\right\rVert_{h}^{\mathbb{R}}\left\lVert[\sigma^{\prime}]\right\rVert_{h}^{\mathbb{R}}.$
On the other hand, using (5.0.7), we have
(5.0.13) $\left\lVert[\Sigma]\right\rVert^{\mathbb{Z}}_{h}\times\left\lVert[\sigma]\right\rVert^{\mathbb{Z}}_{h}$
(5.0.14) $>$ $\lambda\mu_{[\Sigma]}\left\lVert[\Sigma]\right\rVert_{h}^{\mathbb{R}}\times\left\lVert[\sigma]\right\rVert^{\mathbb{R}}_{h}$
(5.0.15) $=$ $\lambda\mu_{[\Sigma]}\nu\left\lVert[\Sigma]\right\rVert_{h}^{\mathbb{R}}\left\lVert[\sigma^{\prime}]\right\rVert^{\mathbb{R}}_{h}.$
Combining the series of inequalities starting from (5.0.8) and (5.0.13), we have
$\left\lVert[\Sigma]\right\rVert^{\mathbb{Z}}_{h}\times\left\lVert[\sigma]\right\rVert^{\mathbb{Z}}_{h}>\lambda\frac{2\mu_{[\Sigma]}}{3v_{g^{\prime}}}\left\lVert[\Sigma]\times[\sigma]\right\rVert^{\mathbb{Z}}_{h}.$
Recall from Definition 1.0.1 that $\mu_{[\Sigma]}$ is a topological invariant independent of metrics, while $v_{g^{\prime}}$ depends only on $g^{\prime}.$ We are done.
6. Remarks about the main theorems
6.1. Other geometric variational problems
Many geometric variational problems can also be formulated on chain spaces over different coefficients, e.g., Almgren-Pitts min-max [40, 27], mean curvature flow [4, 51], Song’s spherical Plateau problem [47], minimizers of elliptic [41] and other general functionals, etc. Motivated by these instances, we propose the following analogue of the Hasse principle:
Conjecture 1.
(The Hasse principle for variational problems) For a variational problem that can be formulated on chain spaces over $\mathbb{R},\mathbb{Z}$ and $\mathbb{Z}/n\mathbb{Z}$ for $n\geq 2,$ the solutions in $\mathbb{Z}$-chains can be reconstructed from the corresponding solutions in $\mathbb{R}$-chains and $\mathbb{Z}/n\mathbb{Z}$-chains for all $n\geq 2$.
For instance, an obvious direction is that of classical Plateau’s problems on Euclidean spaces.
It is worth emphasizing that in number theory, a great deal of research revolves around obstructions to the Hasse principle. For example, the Tate–Shafarevich group is heuristically interpreted as the obstruction to the Hasse principle [28, Sections 7, 8], and it plays a central role in many of the field’s deepest conjectures, such as the Birch and Swinnerton–Dyer conjecture. This naturally leads us to ask:
Conjecture 2.
What are the obstructions to the Hasse principle for variational problems?
6.2. Plateau problems
The analogues of Theorem 2 hold for Plateau problems whose boundaries are compact supported smooth integral currents with finitely many connected components, and the proof is a straightforward adaptation of the one in Chapter 3, using Allard’s monotonicity formula for stationary varifolds with smooth boundaries in [2].
However, once we allow more general boundary chains, the situation becomes much more subtler without Allard’s monotonicity formulas. However, the following conjecture illustrates that considering more general boundary conditions is very natural.
Conjecture 3.
Analogues of Theorem 2 hold for Plateau problems in Euclidean space, when the prescribed boundary is the intersection with the unit sphere of an area-minimizing cone in integral homology .
In other words, we conjecture that an area-minimizing cone in integral homology is an area-minimizing in $\mathbb{Z}/n\mathbb{Z}$ homologies for integers $n$ large enough. Note that intersections with unit spheres of area-minimizing cones in integral homology can be very singular.
6.3. Why real homology is needed
In the statement of Theorem 2, there is no appearance of $H_{d}(M,\mathbb{R}).$ The reader might wonder why we still say the theorem manifests the Hasse principle, which needs both real solutions and solutions modulo prime powers.
Information from $H_{d}(M,\mathbb{R})$ in Section 3.2 is used to give lower bounds on the growth of $\left\lVert\cdot\right\rVert_{g}^{\mathbb{Z}}$ on $H_{d}(M,\mathbb{Z}).$ Heuristically speaking, the Banach space $H_{d}(M,\mathbb{R})$ with norm $\left\lVert\cdot\right\rVert_{g}^{\mathbb{R}}$ represents the asymptotic behavior of $H_{d}(M,\mathbb{Z})$ with norm $\left\lVert\cdot\right\rVert_{g}^{\mathbb{Z}}$ at infinity.
Recall Fact (2.5.3):
$\lim_{\begin{subarray}{c}k\to\infty\\ k\in\mathbb{N}\end{subarray}}\frac{\left\lVert k[\Sigma]\right\rVert_{g}^{\mathbb{Z}}}{k}=\left\lVert[\Sigma]\right\rVert_{g}^{\mathbb{R}}.$
In sharp contrast, by (4.1.2) and (4.1.4), we have constructed metrics $h$ with $\left\lVert[\Sigma]\right\rVert_{h}^{\mathbb{Z}}>\lambda\left\lVert\mu_{[\Sigma]}[\Sigma]\right\rVert_{h}^{\mathbb{Z}}$ for any $\lambda>1.$ The only condition we used about $\mu_{[\Sigma]}$ is that $\mu_{[\Sigma]}[\Sigma]$ admits a smooth representative. By using Lemma 2.1.2, our proof of Lemma 4.1.1 can be modified to show that
Corollary 6.3.1.
If $d\geq 1,c\geq 2,$ $\Sigma]$ is a free class and $M$ is an orientable, and $K$ is the infinite subset of $\mathbb{Z}_{\geq 2}$ defined in Lemma 2.1.2, then for any two integer $k,l\in K$ with $k<l$ and any real numbers $\lambda>1$,  there exist a non-empty open set $\Omega^{\prime}_{[\Sigma],k,l,\lambda}$ in the space of Riemannian metrics, depending on $\Sigma]$, $k,$ $l$, and $\lambda,$ such that for any metric $h\in\Omega^{\prime}_{[\Sigma],k,l,\lambda},$ we have
(6.3.1) $\frac{\left\lVert k[\Sigma]\right\rVert_{h}^{\mathbb{Z}}}{k}>\lambda\frac{\left\lVert l[\Sigma]\right\rVert_{h}^{\mathbb{Z}}}{l}.$
In other words, Federer’s asymptotic (2.5.3) is sharp and cannot be improved to a finitary estimate on the growth of $\left\lVert k[\Sigma]\right\rVert_{h}^{\mathbb{Z}}$ for all $k\in\mathbb{N}$. An alternative way to say this is that the homogeneity of the norm $\left\lVert\cdot\right\rVert_{h}^{\mathbb{Z}}$ fails tremendously. One should also compare with the results in [33, 57, 56], where the authors are considering the failure of homogeneity of $\left\lVert\cdot\right\rVert_{h}^{\mathbb{Z}}$ of Plateau problems in Euclidean space.
In summary, the failure of homogeneity of $\left\lVert\cdot\right\rVert_{g}^{\mathbb{Z}}$ means we cannot control the growth of $\left\lVert\cdot\right\rVert_{g}^{\mathbb{Z}}$ on $H_{d}(M,\mathbb{Z})$ without information from $H_{d}(M,\mathbb{R})$.
6.4. What happens when $\tau_{d}$ does not divide $n$ in Theorem 2
[…]
Corollary 6.4.1.
We have
[…]
Proof.
The denominator of (6.4.1) equals
(6.4.2) $\#\left(H_{d}(M,\mathbb{Z})\right)^{\mathbb{Z}/n\mathbb{Z}}=n^{b_{d}}\prod_{i\in I}\operatorname{gcd}(p_{i}^{\nu_{i}},n).$
Using the notation of Lemma 2.8.1, Subsection 3.1 and Fact 3.2.1 show that whenever
$C_{g}\left\lVert[\Sigma]\right\rVert_{g}^{\mathbb{Z}/n\mathbb{Z}}<\frac{n}{2},$
we have
(6.4.3) $\mathbf{MIN}_{g}^{\mathbb{Z}/n\mathbb{Z}}\left([\Sigma]\right)\subset\bigcup_{[\Pi]\in H_{d}(M,\mathbb{Z})}\left(\mathbf{MIN}_{g}^{\mathbb{Z}}\left([\Pi]\right)\right)^{\mathbb{Z}/n\mathbb{Z}}.$
Expand $\Sigma]$ with respect to the basis in Assumption 2.1.1, we have
(6.4.4) $[\Sigma]=\sum_{j=1}^{b_{d}}a_{j}f_{j}+\sum_{i\in I}b_{i}v_{i},$
for $\bigcup_{j=1}^{b_{d}}\{a_{j}\}\subset\mathbb{N},$ and $\bigcup_{i\in I}\{b_{i}\}\subset\mathbb{Z}\cap(-\frac{n}{2},\frac{n}{2}].$ Apply Lemma 2.5.3 and Fact 2.5.6, we obtain
$\left\lVert[\Sigma]\right\rVert_{g}^{\mathbb{Z}}\leq\upsilon(g)\left\lVert[\Sigma]\right\rVert^{\mathbb{R}}_{g}\leq\upsilon(g)b_{d}\max_{1\leq j\leq b_{d}}|a_{j}|\max_{1\leq j\leq b_{d}}\left\lVert f_{j}\right\rVert^{\mathbb{R}}_{g}.$
This implies whenever
$\max_{1\leq j\leq b_{d}}|a_{j}|<\frac{n}{2C_{g}\upsilon(g)b_{d}\max_{1\leq j\leq b_{d}}\left\lVert f_{j}\right\rVert^{\mathbb{R}}_{g}},$
equation (6.4.3) holds. This implies that the nominator of (6.4.1) is at least
(6.4.5) $\left(\frac{n}{2C_{g}\upsilon(g)b_{d}\max_{1\leq j\leq b_{d}}\left\lVert f_{j}\right\rVert^{\mathbb{R}}_{g}}\right)^{b_{d}}.$
Combining (6.4.2) and (6.4.5), we are done. ∎
Conjecture 4.
We conjecture that the value of the left-hand side of (6.4.1) is $1.$
6.5. Lavrentiev gaps
Another way to think of our main theorems and corollaries from the variational perspective is to view them as the manifestation of Lavrentiev gap phenomena among the norms $\left\lVert\cdot\right\rVert_{g}^{R}$ with $R=\mathbb{Z},\mathbb{R},\mathbb{Z}/n\mathbb{Z}$ for $n\in\mathbb{Z}_{\geq 2}.$
Definition 6.5.1.
Define
(6.5.1) $\mathbf{LavGap}([\Sigma],R)=\{g|\left\lVert[\Sigma]\right\rVert_{g}^{\mathbb{Z}}>\left\lVert[\Sigma]\right\rVert_{g}^{R}\}.$
By Fact 2.5.1, we have
Fact 6.5.1.
$\mathbf{LavGap}([\Sigma],R)$ is always a (possibly empty) open set.
By Fact 2.6.3,  $\mathbf{LavGap}([\Sigma],\mathbb{R})$ has the geometric meaning  that
Fact 6.5.2.
$\mathbf{LavGap}([\Sigma],\mathbb{R})$ is the set of metrics $g$ where no element of $\mathbf{MIN}_{g}^{\mathbb{Z}}([\Sigma])$ can be calibrated.
Theorem 2 and Theorem 3 translate to
  • the infinite intersection $\cap_{n}\mathbf{LavGap}([\Sigma],\mathbb{Z}/n\mathbb{Z})^{\complement}$ for $\tau_{d}|n$ and $n$ large enough, is a closed set with non-empty interior,
  • if $c\geq 2$, $\Sigma]$ is free, and $M$ is orientable, then for $\operatorname{gcd}\left(n,\mu_{[\Sigma]}\right)=1$, the open set $\mathbf{LavGap}([\Sigma],\mathbb{Z}/n\mathbb{Z})$ is non-empty.
  • if $c\geq 2$, $\Sigma]$ is free, and $M$ is orientable, then the open set $\mathbf{LavGap}([\Sigma],\mathbb{R})$ is non-empty.
Using this formulation, one can raise many different problems, for instance
Conjecture 5.
(Camillo De Lellis) The open set $\mathbf{LavGap}([\Sigma],\mathbb{R})$ is dense in the space of smooth Riemannian metrics.
A direct corollary of the above conjecture of De Lellis is that $\mathbb{Z}$-area-minimizing submanifolds are generically not calibrated.
6.6. Naturalness of Theorem 3
We want to remark that $\mathbf{LavGap}([\Sigma],\mathbb{R})$ might appear naturally even when we only consider classical Riemannian manifolds with canonical metrics. For instance, we have
Corollary 6.6.1.
If $\Sigma]\in H_{d}(\mathbb{CP}^{\frac{d+c}{2}},\mathbb{Z}),$ the Fubini-Study metric lies in the closure of $\mathbf{LavGap}([\Sigma],\mathbb{R})$.
Here necessarily both $d,c$ are even and at least $2.$
Proof.
The class $2[\Sigma]$ always admits a connected embedded smooth complex algebraic subvariety $N$ as a representative. For instance, one can take Fermat hypersurfaces in an equator $\mathbb{CP}^{\frac{d}{2}+1}$ of $\mathbb{CP}^{\frac{d+c}{2}}.$ Then by [18, page 70 Example I], $N$ is $\mathbb{Z}$-area-minimizing in the Fubini-Study metric $g_{FS}$ and calibrated by the form $\frac{1}{(d/2)!}\omega^{d},$ where $\omega$ is the standard Kahler form in the Fubini-Study metric. Let $\beta$ be a smooth function on $\mathbb{CP}^{\frac{d}{2}+1}$ such that $\beta\geq 1$ and $\beta^{-1}(1)=N.$ Since $\beta g_{FS}\geq g_{FS}$ as quadratic forms, by the definition of comass in (2.6.1), in metric $\beta g_{FS},$ the form $\frac{1}{(d/2)!}\omega^{d}$ still calibrates $N$. This implies that
(6.6.1) $\left\lVert[\Sigma]\right\rVert_{\beta g_{FS}}^{\mathbb{R}}=\frac{1}{2}\left\lVert 2[\Sigma]\right\rVert_{\beta g_{FS}}^{\mathbb{R}}=\frac{1}{2}\mathbf{M}_{\beta g_{FS}}^{\mathbb{Z}}(N).$
Since $\frac{1}{(d/2)!}\omega^{d}$ have comass $1$ in $\beta g_{FS}$ only on $N,$ by the constancy theorem [36, Section 4.9], any integral current calibrated by $\frac{1}{(d/2)!}\omega^{d}$ in $\beta g_{FS}$ must be an integer multiple of $N$. As $\Sigma]$ is not a torsion class, we deduce that
Fact 6.6.1.
NN is the unique $\mathbb{Z}$-rectifiable current in $2[\Sigma]$ calibrated by $\frac{1}{(d/2)!}\omega^{d}$.
If $\left\lVert[\Sigma]\right\rVert^{\mathbb{R}}_{\beta g_{FS}}=\left\lVert[\Sigma]\right\rVert^{\mathbb{Z}}_{\beta g_{FS}}$, then by (6.6.1), any element $T\in\mathbf{MIN}_{\beta g_{FS}}^{\mathbb{Z}}([\Sigma])$ must satisfy $\mathbf{M}_{\beta g_{FS}}^{\mathbb{Z}}(T)=\frac{1}{2}\mathbf{M}_{\beta g_{FS}}^{\mathbb{Z}}(N).$ Thus, we have $2T\in 2[\Sigma]$ and $2T$ has the same area as $N.$ By Fact 2.6.1, this implies that $2T$ is also calibrated by $\frac{1}{(d/2)!}\omega^{d}$, which contradicts Fact 6.6.1. Replacing $\beta$ by $1+\epsilon(\beta-1)$ for positive $\epsilon\to 0$. We are done.   ∎
Using Almgren’s constructions in [13, Section 5.11], similar results to Corollary 6.6.1 also hold for the standard flat tori $\mathbb{R}^{d+c}/\mathbb{Z}^{d+c}.$
7. Further analogies with number theory
7.1. Ostrowski’s Theorem I: the absolute value norm
Recall the definition of $p$-adic norms  [21, Chapter I]. For readers with experience in number theory, the first question that comes to mind might be why we are not using $R=\mathbb{Z}_{p}$, i.e., the $p$-adic integers, as the Hasse principle is usually stated via the $p$-adic integers to simultaneously capture information from all solutions modulo all prime powers. Roughly speaking, the reason is that we now have both algebraic and analytic structures, and we need both structures to be compatible across different coefficient rings.
By Ostrowski’s theorem [21, Chpater I], up to equivalence, on $\mathbb{Q}$ there are only three types of norms that respect multiplication,
  • the absolute value norm,
  • the $p$-adic norms,
  • the trivial norm.
Here by respecting multiplication, we mean a norm $\left\lVert\cdot\right\rVert$ satisfying $\left\lVert pq\right\rVert=\left\lVert p\right\rVert\left\lVert q\right\rVert$ for all $p,q\in\mathbb{Q}$. The trivial norm is defined to be 0 on the zero element of a ring $R$ and $1$ otherwise.
The author’s heuristic is that for each norm in Ostrowski’s theorem, we have a slightly different version of the Hasse principle.
By Definition 1.0.8, $\mathbb{Z}$-, $\mathbb{R}$- and $\mathbb{Z}/n\mathbb{Z}$-flat chains are defined using the absolute value norm and its induced norms on $\mathbb{Z}/n\mathbb{Z}.$  Our theorems can be interpreted as the Hasse principle with respect to the absolute value norm. There are two more canonical norms by Ostrowski’s theorem.
7.2. Ostrowski’s Theorem II: $p$-adic norms
By [43, Chapter II.1.1], $\mathbb{Z}_{p}$ is the inverse limit of $\mathbb{Z}/p^{n}\mathbb{Z}$ with natural numbers $n\to\infty$ and the system of homomorphisms being mod $p^{n}$ reductions. On one hand, this inverse limit system is not compatible with the norm $\left\lVert\cdot\right\rVert^{\mathbb{Z}/p^{n}\mathbb{Z}}$ in Definition 1.0.8. On the other hand, the $p$-adic norm does naturally induce a norm in the sense of Definition 1.0.8 on $\mathbb{Z}/p^{n}\mathbb{Z}$, i.e., no longer respecting multiplication. The norm is defined by applying the $p$-adic norm to the unique lifts from $\mathbb{Z}/p^{n}\mathbb{Z}$ to $\mathbb{Z}\cap[0,p^{n}-1),$ which we will also call the $p$-adic norms on $\mathbb{Z}/p^{n}\mathbb{Z}$. We conjecture that
Conjecture 6.
(The Hasse principle for $p$-adic variational problems) For a variational problem that can be formulated on chain spaces over $\mathbb{Q}_{p},\mathbb{Z}_{p}$ and $\mathbb{Z}/p^{n}\mathbb{Z}$ for $n\geq 1$, all equipped with the  $p$-adic norm, the solutions in $\mathbb{Z}_{p}$-chains can be reconstructed from the corresponding solutions in $\mathbb{Q}_{p}$-chains and $\mathbb{Z}/p^{n}\mathbb{Z}$-chains for all $n\geq 1$.
Here $\mathbb{Q}_{p}$ is the $p$-adic rational numbers, the fraction field of $\mathbb{Z}_{p}$. The choice of $\mathbb{Q}_{p}$ is to mimic the role of $\mathbb{R}$ as $\mathbb{R}$ is the completion of the fraction field $\mathbb{Q}$ of $\mathbb{Z}$ with respect to the absolute value norm.
7.3. Asymptotic behavior of $\mathbb{Z}/n\mathbb{Z}$-area-minimizing representatives as $n\to\infty$
If we want a limit of $\mathbb{Z}/n\mathbb{Z}$ compatible with $\left\lVert\cdot\right\rVert^{\mathbb{Z}/n\mathbb{Z}}$ as $n\to\infty$, we can renormalize the norm on $\mathbb{Z}/n\mathbb{Z}$ to $\frac{1}{n}\left\lVert\cdot\right\rVert^{\mathbb{Z}/n\mathbb{Z}},$ and take the direct limit of $\mathbb{Z}/n\mathbb{Z}$ with the homomorphisms being $\times m$ multiplications from $\mathbb{Z}/n\mathbb{Z}$ to $\mathbb{Z}/nm\mathbb{Z}.$ The resulting direct limit of $\mathbb{Z}/n\mathbb{Z}$ as $n\to\infty$ is $\mathbb{Q}/\mathbb{Z}$ and the norm is the absolute value norm applied to the unique lifts from $\mathbb{Q}/\mathbb{Z}$ to $\frac{1}{2},\frac{1}{2}).$ Recall Federer’s result (2.5.3) and Section 6.3. We conjecture that
Conjecture 7.
$\mathbb{R}/\mathbb{Z}$-area-minimizing flat chains represents the asymptotic behavior of $\mathbb{Z}/n\mathbb{Z}$-area-minimizing flat chains as $n\to\infty$, just as $\mathbb{R}$-area-minimizing flat chains represents the asymptotic behavior of $\mathbb{Z}$-area-minimizing flat chains.
Here we use $\mathbb{R}/\mathbb{Z}$ instead of $\mathbb{Q}/\mathbb{Z}$ as we want our ring to be complete with respect to our choices of norms.
7.4. Ostrowski’s Theorem III: the trivial norm
Now, the only norm in Ostrowski’s theorem that we have not discussed is the trivial norm. By [53, 55] the trivial norm induces the so-called size, i.e., Hausdorff measure of rectifiable sets, on rectifiable chains. We will follow the literature convention and call area-minimizing flat chains with respect to the trivial norm size-minimizing flat chains.
Unfortunately, $\mathbb{Z}$-flat chains with respect to the trivial norm do not enjoy compactness theorems, so finding size-minimizing representatives in integral homology is a wide-open problem. The only progress so far is by Frank Morgan [37] in some subcases of codimension $c=1$. On the other hand, by [53, 55],  $\mathbb{Z}/n\mathbb{Z}$-size-minimizing flat chains always exist and they enjoy good regularity:
Fact 7.4.1.
A $d$-dimensional $\mathbb{Z}/n\mathbb{Z}$-size-minimizing flat chain $T$ satisfies
  • ([52]) the $d$-th stratum in the Almgren stratification of $T$ is a smooth open not necessarily connected $d$-dimensional submanifold,
  • ([31, 45]) near each point in the (d1)(d-1)-th stratum in the Almgren stratification of $T$, the support of $T$ decomposes into three $d$-dimensional submanifolds with boundaries meeting along a (d1)(d-1)-dimensional submanifold at equal angles of $\frac{2\pi}{3}$, i.e., forming $Y$-shaped triple junctions.
We conjecture that
Conjecture 8.
(The Hasse principle for size-minimizing flat chains) $\mathbb{Z}$-size-minimizing representatives of an integrl homology class $\Sigma]$ exist and can be reconstructed from $\mathbb{R}$-size-minimzing representatives of $\Sigma]^{\mathbb{R}}$ and $\mathbb{Z}/n\mathbb{Z}$-size-minimizing representatives of $\Sigma]^{\mathbb{Z}/n\mathbb{Z}}$ for all $n\in\mathbb{Z}_{\geq 2}$.
7.5. Abelian varieties
There are many more analogies one can make with Diophantine equations. Perhaps the simplest one is that of elliptic curves, or in general, abelian varieties.
Definition 7.5.1.
For a subgroup $G\subset H_{d}(M,R)$, we say that a triple $(G,M,g)$ is $R$-area-abelian, if for any $d$-dimensional $R$-coefficient homology classes $\Sigma],[\sigma],[\sigma^{\prime}]\in G,$ satisfying
$[\Sigma]=[\sigma]+[\sigma^{\prime}],$
and for any element $T\in\mathbf{MIN}^{R}_{g}([\Sigma]),$ there exist $T^{\prime}\in\mathbf{MIN}^{R}_{g}([\sigma]),T^{\prime\prime}\in\mathbf{MIN}^{R}_{g}([\sigma^{\prime}])$ such that
$T=T^{\prime}+T^{\prime\prime}.$
An $R$-area-abelian triple $(G,M,g)$ is a natural analogue of abelian varieties in the coefficient ring $R.$
The simplest example of an area-abelian triple is that of $G\cong\mathbb{Z},R=\mathbb{Z}$ and the generator $G$ admitting a calibrated $\mathbb{Z}$-area-minimizing representative. As a corollary of [59, Theorem 3.19, 3.20], when the codimension $c$ satisfies $c\geq 3,$ and $H_{d}(M,\mathbb{Z})$ is torsion-free, we can find metrics $g$ on $M$ such that for some finite index subgroup $G$ of $H_{d}(M,\mathbb{Z})$, $(G,M,g)$ is a $\mathbb{Z}$-area-abelian triple and $(H_{d}(M,\mathbb{R}),M,g)$ is an $\mathbb{R}$-area-minimizing triple. Thus, area-abelian triples are abundant. A nature question is
Conjecture 9.
What are the characterizing properties of area-abelian triples?
We can also formulate the following analogue of Birch and Swinnerton–Dyer conjecture
Conjecture 10.
For a $\mathbb{Z}$-area-abelian triple $(G,M,g)$, the rank of $G$ can be recovered from the collection of cardinalities of all subgroups $H\subseteq G^{\mathbb{Z}/n\mathbb{Z}}$ such that $(H,M,g)$ forms a $\mathbb{Z}/n\mathbb{Z}$-area-abelian triple, for every $n\in\mathbb{Z}_{\geq 2}$.
One can also interpret the above conjecture as follows: when $\mathbb{Z}$-area-abelian subgroups $G$ exist, a stronger form of Theorem 2 holds uniformly across all elements of $G.$
7.6. Rational solutions
In number theory, one often considers rational rather than integral solutions to Diophantine equations, since the field structure of $\mathbb{Q}$ allows the use of many additional algebraic theories.
In the setting of solving for area-minimizing representatives, integral homology classes always admit $\mathbb{Z}$-area minimizing representatives, in contrast with the case of Diophantine equations. However, the situation is subtler for rational homology: a rational homology class does not necessarily admit a $\mathbb{Q}$-area-minimizing representative, as $\mathbb{Q}$ is not complete with respect to the absolute value norm. From this perspective, a more faithful analogy with Diophantine equations is to study the existence of $\mathbb{Q}$-area-minimizing representatives of rational homology classes.
By [53], whenever they exist, $\mathbb{Q}$-area-minimizing representatives are rectifiable and therefore satisfy monotonicity formulas. Moreover, the field structure of $\mathbb{Q}$ ensures that $\|\cdot\|^{\mathbb{Q}}$ defines a norm on $H_{d}(M,\mathbb{Q})$ satisfying homogeneity. Motivated by this, we propose the following conjecture:
Conjecture 11.
A rational homology class $\Pi]\in H_{d}(M,\mathbb{Q})$ admits a $\mathbb{Q}$-area-minimizing representative in a metric $g$ if and only if the real reduction $\Pi]^{\mathbb{R}}$ admits an $\mathbb{R}$-area-minimizing representative in the metric $g$ that can be expressed as a countable $\mathbb{Q}$-linear combination of $\mathbb{Z}$-area-minimizing representatives in the metric $g$.
Again, as a corollary of [59, Theorem 3.19, 3.20], Riemannian manifolds admitting $\mathbb{Q}$-area-minimizing representative in all rational homology classes of codimension at least $3$ are abundant. Based on these examples, we further conjecture that
Conjecture 12.
A rational homology class $\Pi]\in H_{d}(M,\mathbb{Q})$ admits a $\mathbb{Q}$-area-minimizing representative in a metric $g$ if and only if $\Pi]$ belongs to a subgroup $G\subset H_{d}(M,\mathbb{Q})$ with $(G,M,g)$ a $\mathbb{Q}$-area-abelian triple.
In other words, we conjecture that $\mathbb{Q}$-area-minimizing representatives are precisely elements of $\mathbb{Q}$-area-abelian subgroups $G$, and thus our Definition 7.5.1 is a canonical analogue of abelian varieties over $\mathbb{Q}.$